### v1

# LocalSearchBench: Benchmarking Agentic Search in Real-World Local Life Services

Abstract.
Recent advances in large reasoning models (LRMs) have enabled agentic search systems to perform complex multi-step reasoning across multiple sources. However, most studies focus on general information retrieval and rarely explores vertical domains with unique challenges. In this work, we focus on local life services and introduce LocalSearchBench, which encompass diverse and complex business scenarios. Real-world queries in this domain are often ambiguous and require multi-hop reasoning across merchants and products, remaining challenging and not fully addressed. As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench includes over 150,000 high-quality entries from various cities and business types. We construct 300 multi-hop QA tasks based on real user queries, challenging agents to understand questions and retrieve information in multiple steps. We also developed LocalPlayground, a unified environment integrating multiple tools for agent interaction. Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.1) achieves only 34.34% correctness, and most models have issues with completeness (average 77.33%) and faithfulness (average 61.99%). This highlights the need for specialized benchmarks and domain-specific agent training in local life services. Code, Benchmark, and Leaderboard are available at localsearchbench.github.io
Agentic Search, Large Reasoning Model, Local life services
*Project leader.
[…]
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2026/06copyright: noneccs: Computing methodologies Natural language processingccs: Computing methodologies Artificial Intellegence
1. Introduction
(a) General multi-hop QA
(b) Local life service single-hop QA
(c) Local life service multi-hop QA
Figure 1. Illustration of different types of QA tasks
Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (xu2025towards; ferrag2025llm; wu2025agentic; li2025webthinker; ke2025survey; chai2025rlfactory). This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (sun2025simpledeepsearch; zhang2025evolvesearch; shi2025pangudeepdiveradaptivesearch; shi2025toollearningwildempowering; zhang2025web; li2025search; mo2025conversational; yue2025promoting; yue2025uiorchestra). These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (zhou2024webarenarealistic; yao2023webshopscalable). However, there has been little work applying agentic search to the local life service domain, and the lack of suitable evaluation benchmarks further constrains new research in this area.
Local life services cover a wide range of sub-domains, including dining, lifestyle, shopping, accommodation, travel, and healthcare (lan2025benchmarking; 10.1145/3580305.3599874). Leveraging location-based digital platforms, these services connect users with nearby merchants and service providers, forming complex networks of businesses, user preferences, and location-dependent service offerings, and creating a vast market (10.1145/3534678.3539122; kusk2025flexible; liu2025mrgrp). As shown in Figure 1(1(c)), real-world user queries often involve multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling, which require complex multi-step processes to deliver the final response. lan2025benchmarking target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios. Therefore, there is an urgent need for a multi-hop benchmark tailored to local life scenarios, accurately reflecting the complexity of real-world search tasks and fully challenging existing agentic search methods.
In this paper, we propose an agentic search framework LocalPlayground to evaluate LRMs for local life service. Given the absence of appropriate benchmarks and the critical need for comprehensive evaluation systems, our approach addresses these challenges through comprehensive dataset construction, systematic evaluation framework and targeted model evaluation.
We introduce LocalSearchBench, a comprehensive benchmark comprising over 150,000 data entries across 6 categories: dining, shopping, accommodation, travel, and healthcare, spanning 3 major cities in China. These entries are transformed from raw data into high-quality augmented and anonymized datasets.
Based on these data, we propose a high-quality instruction synthesis method to design 300 multi-hop QA tasks across three reasoning dimensions: multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling. These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
We build a comprehensive evaluation environment with RAG based on LocalSearchBench data and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
In summary, our contributions are threefold.
  • To the best of our knowledge, we are the first to build a comprehensive offline high-quality real-world agentic search benchmark for local life service.
  • We propose a systematic agentic search framework LocalPlayground and tool environment for evaluating and applying LRMs in local life services.
  • We evaluate existing models and find that untrained models struggle to complete these tasks, while trained models perform poorly on domain-specific professional requirements.
2. Related Work
2.1. Agentic Search
gou2025mind2web2evaluatingagentic define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks. Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (yang2025agentic). They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (li2025towards).
Recent advances in the reasoning abilities of LRMs further foster the development of LRM-based agentic search systems. The most representative products include OpenAI Deep Research (jaech2024openai), Gemini Deep Research (comanici2025gemini), and Tongyi Deep Research (tongyidr). In addition to accessing real-time information through search APIs, these systems leverage various auxiliary tools to enable deeper and broader research on highly complex problems. In the open-source community,  li2025search first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.  li2025webthinker,  jin2025search, and  song2025r1 further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
These methods have made significant progress in general information retrieval and question answering. However, there has been little advancement in vertical domains, especially in areas related to local lifestyle services, due to the challenges associated with training domain-specific agents and the lack of high-quality evaluation benchmarks. In this paper, we propose LocalSearchBench, a comprehensive benchmark specifically designed for evaluating LRMs in local life service scenarios.
2.2. Benchmarking Agentic capabilities
Existing benchmarks for agentic search systems in general domain can be broadly categorized into closed-ended and open-ended QA tasks. Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (yang2018hotpotqa; ho2020constructing; press2022measuring; trivedi2022musique), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (wei2025browsecomps; zhou2025browsecompzhbenchmarking; xi2025infodeepseek; pham2025sealqa), and fact-checking tasks that evaluate claim verification abilities (wei2024long; wadden2020fact; jiang2020hover; wang2024mfc). Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (rosset2024researchy) and expert-level research tasks (du2025deepresearch; bosse2025deep; tan2024proxyqa), with some incorporating multi-modal capabilities (yuan2025videodeepresearch).
Despite covering various reasoning abilities, existing benchmarks differ from real user queries in local life services and lack assessment of multi-domain coordination. Users often face complex tasks requiring simultaneous service bookings and constraint satisfaction, which current datasets do not adequately evaluate.
However, although these benchmarks cover a range of reasoning capabilities, they have significantly different distribution from real user queries in local life service scenarios, lacking assessment of the multi-domain service coordination required in real-world applications. In reality, users often need to accomplish complex tasks that require coordinating multiple services, such as simultaneously booking restaurants, arranging transportation, and planning entertainment, all while satisfying geographic and temporal constraints. Existing specialized datasets are insufficient for evaluating challenging scenarios like event planning that require multiple API calls across different service domains and real-time constraint satisfaction.
We propose LocalSearchBench to assess multi-hop reasoning and coordinated retrieval for complex local life service tasks, and LocalPlayground as evaluation frameworks that comprehensively measure performance in these specialized scenarios.
3. Construction of LocalSearchBench
Figure 2. Workflow of Local Merchant Database Construction
In this section, we introduce LocalSearchBench, a novel benchmark for agentic search in local life services. LocalSearchBench consists of two main components: Local Merchant Database (Section 3.1) and a multi-hop QA benchmark (Section 3.3).
3.1. Local Merchant Database Construction
In this subsection, we present the construction of the local merchant database for LocalSearchBench, which serves as the foundation for multi-hop QA tasks. As shown in Figure 2, our database construction consists of two stages: Merchant Seed Data Collection and Data Augmentation and Synthesis.
3.1.1. Merchant Seed Data Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we built our dataset based on source data from a leading local life service platform M. We sampled over 260,000 real merchant records spanning from January 2025 to December 2025, covering diverse service categories and geographical distributions across three major Chinese cities, with the generation of these seed data requiring 114 hours of computation time.
Multi-Scenario Coverage
Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 11. We curated a keyword pool of 411 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
Multi-City Coverage
The collection spans 3 major Chinese cities with comprehensive coverage of benchmarks (Table 1). The location database includes 859 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas. Figure 3 shows word clouds of landmark distributions across the 3 cities.
Table 1. Geographic coverage across 3 major Chinese cities
City Landmarks Districts
Shanghai 295 16
Beijing 287 16
Guangzhou 277 11
Total 859 43
(a) Shanghai
(b) Beijing
(c) Guangzhou
Figure 3. Word clouds showing the landmark locations across 3 major Chinese cities.
Merchant Seed Data Retrieval Query Generation
The system employed scenario-specific keyword pools, constructing queries through random combinations of 1-2 keywords to simulate natural search patterns. This approach generated diverse query variations from the 411-keyword foundation, ensuring comprehensive coverage of local search intents across regional preferences.
This multi-city keyword strategy captured regional variations across 43 administrative divisions, providing robust coverage of China’s diverse local search ecosystems as the foundation for LocalSearchBench database synthesis.
3.1.2. Data Augmentation and Synthesis
Data Augmentation
The original merchant seed data contains only 12 basic fields with many missing or unstandardized attributes, which is inadequate for agentic search tasks that require dense, high-coverage merchant profiles. We employ a Data Augmentation Agent to expand these 12-field records into 29-field profiles using LLMs across six dimensions (see Figure 2 and Figure 12). The agent enriches each merchant’s basic, locational, operational, service, contact, and promotional information through 17 newly added fields, as illustrated in Figure 2. The enhanced dataset covers: (1) basic information with standardized categories and descriptions; (2) location intelligence including structured geographical data, coordinates, and landmarks; (3) business operations such as operating hours and pricing; (4) service characteristics like facility lists; (5) contact information, including phone numbers and emails; and (6) promotional content, such as product descriptions and discount offers.
Privacy Rewriting
To comply with data protection regulations and enable safe public release, we employ a Privacy Rewriting Agent on the augmented data (see Figure 13). Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized. The agent rewrites 6 privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
3.1.3. Data Validation
A Validation Agent, implemented as an LLM-as-a-Judge framework, conducts a comprehensive, two-stage quality assessment to ensure both data quality and privacy compliance in the generated database. Detailed instructions have been carefully designed for augmented data validation and privacy rewriting validation, as illustrated in Figure 14.
Augmented Data Validation
It evaluates augmented Data quality against original seed data across six core dimensions: (1) Data Completeness (20% weight) assessing field coverage and information richness; (2) Logical Consistency (30% weight) validating business logic coherence and field correlations including name format compliance, category alignment, and pricing rationality; (3) Geographic Accuracy (10% weight) verifying location information correctness and address format consistency; (4) Content Quality (20% weight) evaluating textual descriptions, language expression, and business relevance; (5) Format Standardization (10% weight) ensuring data format compliance and structural consistency; and (6) Innovation Degree (10% weight) measuring content creativity and merchant name diversity through batch-level uniqueness analysis. Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
Privacy Rewriting Data Validation
It evaluates privacy rewriting data quality against augmented data across four core dimensions: (1) Anonymization Degree (50% weight) ensuring complete merchant identity anonymization including name deidentification, location sanitization, and contact information anonymization; (2) Logical Consistency (20% weight) maintaining field relationships and business logic coherence after privacy rewriting; (3) Content Consistency (20% weight) preserving non-sensitive information alignment with previous stage outputs; and (4) Innovation Degree (10% weight) assessing the creativity and diversity of anonymized merchant names to prevent pattern recognition. Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
Table 2. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Evaluation Dimension Weight Evaluation Fields
Data Completeness 20% All
Logical Consistency 30% name, category, subcategory, price_range, business_hours, facilities
Geographic Accuracy 10% city, district, landmark, address
Content Quality 20% description, rating, review_count, tags, promotions, products, group_deals
Format Standardization 10% phone, mobile, email
Innovation Degree 10% name
(b) Privacy Writing Data Quality
Evaluation Dimension Weight Evaluation Fields
Anonymization Degree 50% name, address, email
Logical Consistency 20% description, phone, mobile
Content Consistency 20% category, subcategory, city, district, landmark, price_range, business_hours, rating, review_count, delivery_available, delivery_fee, delivery_range, min_order_amount
Innovation Degree 10% name
3.1.4. Local Merchant Database Statistics
Overall, we curated 150,031 representative samples from a pool of over 260,000 original seed records to create synthesized data, which constitute our LocalSearchBench dataset.
The proportions of each category, the number of data for each city are presented in Figure 4. The distribution closely mirrors that of the source data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms. The geographical distribution of merchant data for each city was derived through geospatial inference to generate latitude-longitude heatmaps, as shown in Figure 5. The distribution patterns align well with urban clustering phenomena and regional economic hierarchies, demonstrating the dataset’s realistic geographical representation.
(a) City distribution
(b) Shanghai
(c) Beijing
(d) Guangzhou
Figure 4. City distribution and category distribution of 150,031 merchant data across three cities.
(a) Shanghai
(b) Beijing
(c) Guangzhou
Figure 5. Geographical distribution heatmaps of 150,031 merchant data across three cities.
3.2. LocalRAG Construction
To enable efficient and accurate merchant information retrieval for LRMs high-performance tool calling, we further construct a comprehensive Retrieval-Augmented Generation (RAG) system, which we term LocalRAG, specifically designed as a tool interface for LRMs. Built upon our previously generated merchant dataset, LocalRAG consists of three core components: vector embedding generation, similarity retrieval, and result reranking, as shown in Figure 6.
3.2.1. Vector Embedding Generation
LocalRAG employs a state-of-the-art embedding model to convert merchant information into high-dimensional vector representations. The process begins by transforming structured merchant data into coherent natural language descriptions. Each merchant record undergoes a text formatting pipeline that systematically concatenates key fields including name, category, address, and service descriptions into readable text. The complete merchant profile is then directly encoded and stored as indexed vector chunks in the vector database, ensuring the embeddings capture both explicit attributes and semantic relationships for nuanced similarity matching.
3.2.2. Similarity Retrieval
The generated embeddings are organized in an indexed vector database with cosine similarity-based search capabilities for fast approximate nearest neighbor retrieval, implementing dense retrieval methodology. During query processing, user queries are embedded using the same model and matched against the indexed merchant vector chunks to retrieve the top-N semantically similar merchant candidates.
3.2.3. Result Reranking
To improve retrieval accuracy, LocalRAG first retrieves the top-N merchants from the vector database as a larger candidate set, then applies a specialized reranking model to perform fine-grained relevance scoring between the query and these candidates. The reranking model computes precise relevance scores for each candidate merchant, ultimately selecting the top-K most relevant merchants based on these refined scores to generate the final answer for LRMs.
Figure 6. Workflow of LocalRAG
3.3. Multi-hop QA Construction
Figure 7. Workflow of Multi-hop QA Construction.
In this subsection, we present the multi-hop QA construction of LocalSearchBench. Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and data validation. Each of these stages is detailed in the following subsubsections.
3.3.1. Seed Question Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we conducted a statistical analysis to understand user search behaviors and service interactions. Based on requirement complexity and reasoning demands, we classify questions into five intelligence levels (Table 3): L1 (Basic) for simple factual queries; L2 (Advanced) for ambiguous or multi-condition searches; L3 (Intelligent) for composite requirements with multi-turn reasoning; L4 (Expert) for personalized planning; L5 (AGI) for cross-platform coordination with adaptive reflection.
Table 3. Intelligence Level Classification for Seed Questions.
Intellig- ence Level Grade Requirement Understanding Planning-Search-Reflection Loop
L1 Basic
Standard expression of single
precise requirement
Supply matching search with $\leq$
2 conditions
L2 Advanced
Ambiguous expression of single
precise requirement;
Standard expression of
single ambiguous requirement
Supply matching search with $\geq$
3 conditions;
Semantic retrieval search supply;
Multi-source information
integration supply
L3 Intelligent
Composite requirements;
Multi-turn conversational
requirements
Reasoning with certain complexity;
Semantic retrieval with external
information dependency;
Procedural planning requirements;
Comparison and contrast scenarios
L4 Expert
Personalized requirements
External hard-to-access dependency;
Open-ended planning requirements
L5 AGI
Cross-platform coordination;
Complex decision-making;
Exception handling and reflection
Cross-platform dependency;
Task execution;
Adaptive reflection
We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across six service categories: dining, lifestyle, shopping, accommodation, healthcare, and tourism. The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs. Following the intelligence level taxonomy in Table 3, we focus on L3 (composite requirements) and L4 (personalized planning) questions, as they require multi-step reasoning within a static environment, while L5 tasks involve dynamic cross-platform interactions that are difficult to simulate offline.
(a) Intelligence Level Distribution
(b) Example Questions
Figure 8. Intelligence Level Analysis for Collected Questions
3.3.2. Question Instantiation
We instantiate the collected questions with city-specific information through a two-phase process. First, we construct city profiles for each of the 3 cities (Section 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc. The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by native annotators to ensure factual accuracy and cultural authenticity.
Second, annotators from platform M manually instantiate questions into concrete multi-hop queries. For each city, approximately 100 questions are selected from the collected 1,200 questions and grounded with city-specific information. To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments (2 for 3-hop, 3 for 4-hop, 4 for 5-hop) based on business relevance and logical coherence. These segments include: cross-merchant comparison, schedule planning, chained events, and constraint resolution. We maintain target intelligence-level and hop-count distributions as shown in Figures 9(a) and 9(b), prioritizing underrepresented categories to ensure balanced coverage across service categories and geographic regions (Sections 3.1.1 and 3.1.1).
(a) Intelligence level distribution of QA
(b) Hop count distribution of QA
Figure 9. Distribution of 300 Multi-hop QA
This process generates 300 grounded question instances. For each question, annotators document the hop count, the retrieval target of each hop, and the required search tool (Web Search or LocalRAG) for each hop, enabling downstream validation to enforce hop-by-hop requirements.
3.3.3. Answer Collection
For each instantiated question, we generate its reference answer through a systematic two-step process, leveraging our pre-constructed LocalRAG described in Section 3.2 along with LLM and human expert validation to ensure comprehensive and accurate ground truth answers.
Step 1: Merchant Retrieval. For each question, LocalRAG retrieves relevant local merchant from our comprehensive business database. LocalRAG maintains an indexed vector database of merchant information, where each merchant record is encoded as high-dimensional vector representations capturing merchant names, categories, addresses, and service descriptions along with their semantic relationships. Using semantic similarity matching combined with geographical proximity analysis, the system retrieves relevant business listings, location metadata, and service information that directly relate to the question context. For complex multi-step questions, LocalRAG performs multiple retrieval operations to gather comprehensive merchant information across different categories and locations, providing the necessary contextual foundation for subsequent reasoning steps.
Step 2: Initial Answer Generation. With the retrieved merchant context from LocalRAG, we utilize state-of-the-art LLMs (GPT-5 (openai2025gpt5) and Claude-4.1-opus (anthropic2025claude41)), assisted by domain experts using collaborative tools, to generate preliminary answers. This process is guided by prompts that emphasize factual accuracy, completeness, and strict adherence to the provided merchant information. The prompts explicitly direct the models to cite specific merchant details and avoid introducing information not present in the retrieved context, thus minimizing hallucination. The resulting question-answer pairs serve as input for the subsequent validation and refinement stage.
3.3.4. Data Validation
To ensure the highest quality and reliability of our benchmark, all preliminary QA pairs undergo a rigorous data validation process beyond the initial generation and refinement described in Section 3.3.3. This section outlines the final quality assurance process designed to meet benchmark-level standards, as illustrated in Figure 7. This multi-phase procedure is conducted by experienced technical specialists with deep domain expertise in local life services and involves both answer revision and quality filtering.
Step 1: Expert Answer Revision. Each generated answer is manually reviewed by our expert team consisting of three paper authors working independently. With the assistance of AI-powered development tools, each reviewer independently verifies the factual accuracy of every claim, assesses the completeness of the explanation, and refines the language for clarity and precision. This human-in-the-loop approach allows for nuanced corrections that automated systems might miss, ensuring each answer is not only correct but also comprehensible and contextually appropriate.
Step 2: Quality Filtering. After revision, the QA pairs are subjected to a final filtering step. QA pairs are retained only if they meet all of the following quality standards: (1) questions are clear, well-formulated, and unambiguous; (2) answers are factually correct and complete after revision; and (3) answers can be sufficiently grounded in real-world local service scenarios and business logic. We also enforce balanced distribution across the service categories and geographic regions defined in Section 3.1.1. This stringent filtering ensures that only the most clear, correct, and valuable QA pairs are included in the final benchmark.
3.3.5. Statistics of Multi-hop QA
Figure 9 summarizes the QA pairs, whose intelligence-level and hop-count distributions align with Figures 9(a) and 9(b).
4. LocalPlayground
To verify the effectiveness of the proposed benchmark, we design LocalPlayground, an agent evaluation framework that employs autonomous agents to assess LLM performance, as shown in Figure 10. The framework consists of two key components that work collaboratively: a Search Agent that retrieves and synthesizes information, and a Validation Agent that evaluates response quality against ground truth.
Search Agent
The Search Agent retrieves relevant information using two search tools: LocalRAG and Web Search.
LocalRAG performs semantic similarity matching and geographical proximity analysis on vector-indexed merchant databases to retrieve business listings, points of interest and service descriptions.
Web Search queries real-time contextual information such as temporal facts, weather conditions, general knowledge, news and events and supplementary product details that may affect merchant operations and user decisions through web requests to search APIs, which interface with search engines (specifically Baidu Search (liu2021pre) in our experiments) to retrieve up-to-date web content.
The Search Agent operates iteratively over N rounds. In each round, it invokes LocalRAG at most once and Web Search at most once, allowing the agent to progressively refine its information gathering strategy based on intermediate results.
Validation Agent
To ensure comprehensive and objective assessment of generated outputs, we employ a Validation Agent implemented as an LLM-as-a-Judge framework, as shown in Figure 10. Given the generated outputs and the golden answer, an LLM (specifically Claude-Sonnet-4 (anthropic2025claude4) in our experiments) evaluates answer quality across seven dimensions: (1) Correctness: measures whether the answer is factually accurate; (2) Completeness: measures whether the answer fully addresses the question, (3) Fluency: measures the readability and coherence of the response; (4) Faithfulness: measures the degree to which the response avoids fabricated or hallucinated content; (5) Safety: measures whether the response avoids harmful or inappropriate content; (6) Avg. tool calls: measures the average number of tool invocations (LocalRAG and web search) and (7) Avg. rounds: measures the average number of conversation rounds.
Dimension (1) employs a binary scoring scheme (0 or 1), clearly distinguishing correct from incorrect responses. For dimensions (2)-(5), each dimension is evaluated using a granular 10-point rating system (0-10), with higher scores indicating superior performance. Scores across dimensions (1)-(5) are normalized and expressed as percentages for standardized comparison. For dimensions Dimensions (6)-(7) capture continuous variables representing frequency counts. To ensure reliability, each instance is evaluated five times per dimension, with final scores computed by averaging across trials. Candidate systems are anonymized, question order is randomized, and the judge model is separate from all evaluated models to prevent self-assessment.
Figure 10. Overview of LocalPlayground
5. Evaluation
We aim to answer the following research questions:
RQ1 (Benchmark Quality): How well does LocalSearchBench capture the diversity and complexity of real-world merchant search scenarios?
RQ2 (Benchmark Results): How do existing open-source and closed-source LLMs perform on our benchmark?
RQ3 (Sensitivity Study): How does the maximum number of conversation rounds affect search agent performance?
5.1. Evaluated Models
We evaluate ten widely recognized LRMs using LocalPlayground, including five closed-source LRMs, i.e., GPT-4.1 (openai2025gpt4-1), Gemini-2.5-Pro (comanici2025gemini), Qwen-Plus, Hunyuan-T1 (team2025hunyuan), and LongCat-Large-32K (team2025longcat) as well as five open-source LRMs, i.e., DeepSeek-V3.1 (deepseek2025v3), GLM-4.5 (zeng2025glm), Qwen3-235B-A22B, Qwen3-32B, and Qwen3-14B (yang2025qwen3).
Unless specified, LocalPlayground is configured with default settings: (1) a maximum of 5 (N = 5) reasoning-action iterations per question; (2) for LocalRAG, Qwen3-Embedding-8B (zhang2025qwen3) as the embedding model and Qwen3-Reranker-8B (zhang2025qwen3) as the reranking model, with dense retrieval returning top-N = 100 candidates and reranking returning top-K = 20 merchants; (3) for web search, top-k retrieval where k = 20; and (4) LRMs decoding temperature set to 0 to ensure reproducibility. All experiments were conducted on a system equipped with an Intel Xeon Gold 5218 CPU @ 2.30GHz and 1 $\times$ H20-141G GPU.
5.2. Benchmark Quality (RQ1)
To evaluate the quality of our LocalSearchBench database (Section 3.1.1), we employ the two-stage validation framework described in Section 3.1.3. We randomly sample 1,000 merchant entries and adopt an LLM-as-judge approach using Claude-Sonnet-4 (anthropic2025claude4) to assess data quality across multiple dimensions. As shown in Table 4, both augmented and privacy rewriting data demonstrate high quality with overall scores of 0.8596 and 0.9217, respectively.
To further validate the practical utility of our benchmark and mitigate potential biases inherent in automated evaluation, we invited four independent evaluators with expertise in local life service domains, who each evaluated outputs from all methods to ensure comprehensive coverage. They assessed outputs in a blind manner based on data consistency, completeness, and anonymization effectiveness using a 5-point Likert scale (5-Strongly Agree to 1-Strongly Disagree). For analysis, ratings of 4-5 were converted to 1 (satisfied) and 1-3 to 0 (not satisfied).
As shown in Table 7, the human evaluation achieves 86.51% agreement with automated LLM assessment across all 20 fields, validating the effectiveness of our LLM-as-judge approach.
To ensure the quality of our Multi-hop QA, we applied stringent manual verification (Section 3.3.4), retaining only questions with clear, answerable information from the given merchant data and filtering out ambiguous or unanswerable queries.
Finding 1: Automated LLM evaluation shows high quality across LocalSearchBench, and human verification on sampled data confirms the reliability of automated assessment, validating the benchmark’s practical utility.
Table 4. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Dimension Weight Score
Data Completeness 20% 0.8605
Logical Consistency 30% 0.9000
Geographic Accuracy 10% 0.8568
Content Quality 20% 0.8521
Format Standardization 10% 0.9060
Innovation Level 10% 0.7080
Overall 100% 0.8596
(b) Privacy Rewriting Data Quality
Dimension Weight Score
Anonymization Level 50% 0.9473
Logical Consistency 20% 0.8383
Content Consistency 20% 0.9919
Innovation Level 10% 0.8200
Overall 100% 0.9217
5.3. Benchmark Results (RQ2)
Impact of web search.
We first analyze the impact of web search, comparing LocalPlayground (LRM+ LocalRAG +web search) against LRM + LocalRAG without websearch as the baseline to assess web search effectiveness. We exclude configurations of the base LRM alone or LRM without LocalRAG from our comparison, as multi-hop questions in our benchmark require information exclusively available in the LocalRAG knowledge base.
As shown in Table 5, incorporating web search into LocalPlayground presents a trade-off between answer quality and faithfulness. On average, LocalPlayground improves Correctness by 4.37pp (from 25.58% to 29.95%) and Completeness by 3.95pp (from 73.38% to 77.33%), but reduces Faithfulness by 3.64pp (from 65.63% to 61.99%). This trade-off is most pronounced for DeepSeek-V3.1, which gains 4.57pp in Correctness but loses 8.45pp in Faithfulness. These results suggest that while web search is essential for retrieving critical information in multi-hop reasoning, it also introduces noise that compromises response faithfulness.
Performance across language models.
The choice of underlying LRM significantly influences performance across all metrics. Among all models, DeepSeek-V3.1 achieves the highest Correctness (34.34%), followed by GLM-4.5 (33.78%) and LongChat-Large-32K (33.19%). Qwen-Plus-Latest leads in Completeness (80.94%), while Gemini-2.5-Pro excels in Faithfulness (78.26%). Closed-source models outperform open-source ones in completeness (79.08% vs. 75.57%) and efficiency (2.25 vs. 2.61 tool calls), though differences in Correctness are more modest. Despite these variations, overall performance remains limited, as even the best model achieves only 34.34% Correctness, underscoring the substantial challenge of multi-hop local search.
Table 5. Performance comparison of different models on LocalSearchBench (Max Round N = 5).
Model Config Evaluation Metrics
Avg. tool calls Avg. rounds Correctness (%) Completeness (%) Fluency (%) Faithfulness (%) Safety (%)
Closed-source LRMs
[Uncaptioned image] GPT-4.1 w/o web search 1.05 2.06 18.73 70.64 72.37 75.10 83.71
LocalPlayground 1.72 ($\uparrow$0.67) 2.70 ($\uparrow$0.64) 26.76 ($\uparrow$8.03) 75.42 ($\uparrow$4.78) 71.54 ($\downarrow$0.83) 72.63 ($\downarrow$2.47) 81.00 ($\downarrow$2.71)
[Uncaptioned image] Gemini-2.5-Pro w/o web search 1.32 2.29 22.56 70.51 74.48 75.52 84.51
LocalPlayground 1.89 ($\uparrow$0.57) 2.86 ($\uparrow$0.57) 26.09 ($\uparrow$3.53) 77.93 ($\uparrow$7.42) 73.50 ($\downarrow$0.98) 78.26 ($\uparrow$2.74) 83.81 ($\downarrow$0.70)
[Uncaptioned image] Qwen-Plus-Latest w/o web search 1.79 2.79 27.76 77.19 73.21 68.78 83.41
LocalPlayground 2.59 ($\uparrow$0.80) 3.12 ($\uparrow$0.33) 32.79 ($\uparrow$5.03) 80.94 ($\uparrow$3.75) 73.65 ($\uparrow$0.44) 68.68 ($\downarrow$0.10) 83.14 ($\downarrow$0.27)
[Uncaptioned image] LongCat-Large-32K w/o web search 1.59 2.49 30.93 79.45 71.82 65.58 80.72
LocalPlayground 2.73 ($\uparrow$1.14) 3.22 ($\uparrow$0.73) 33.19 ($\uparrow$2.26) 80.51 ($\uparrow$1.06) 73.40 ($\uparrow$1.58) 60.80 ($\downarrow$4.78) 81.14 ($\uparrow$0.42)
[Uncaptioned image] Hunyuan-T1 w/o web search 1.82 2.65 30.10 74.50 73.53 71.76 84.50
LocalPlayground 2.30 ($\uparrow$0.48) 3.15 ($\uparrow$0.50) 32.94 ($\uparrow$2.84) 80.59 ($\uparrow$6.09) 73.00 ($\downarrow$0.53) 63.77 ($\downarrow$7.99) 84.71 ($\uparrow$0.21)
Open-source LRMs
[Uncaptioned image] Qwen3-14B w/o web search 1.20 2.16 21.07 71.54 70.67 54.47 82.71
LocalPlayground 2.04 ($\uparrow$0.84) 2.85 ($\uparrow$0.69) 26.42 ($\uparrow$5.35) 76.25 ($\uparrow$4.71) 70.97 ($\uparrow$0.30) 48.54 ($\downarrow$5.93) 81.54 ($\downarrow$1.17)
[Uncaptioned image] Qwen3-32B w/o web search 1.33 2.10 22.74 68.23 70.40 47.73 80.80
LocalPlayground 2.48 ($\uparrow$1.15) 2.89 ($\uparrow$0.79) 24.75 ($\uparrow$2.01) 71.37 ($\uparrow$3.14) 70.17 ($\downarrow$0.23) 40.84 ($\downarrow$6.89) 81.34 ($\uparrow$0.54)
[Uncaptioned image] Qwen3-235B-A22B w/o web search 1.26 2.10 23.08 66.56 70.50 54.94 81.20
LocalPlayground 2.35 ($\uparrow$1.09) 2.81 ($\uparrow$0.71) 28.43 ($\uparrow$5.35) 73.48 ($\uparrow$6.92) 70.54 ($\uparrow$0.04) 52.43 ($\downarrow$2.51) 81.67 ($\uparrow$0.47)
[Uncaptioned image] GLM-4.5 w/o web search 1.92 2.92 29.10 76.56 67.86 73.15 80.97
LocalPlayground 2.73 ($\uparrow$0.81) 3.66 ($\uparrow$0.74) 33.78 ($\uparrow$4.68) 76.76 ($\uparrow$0.20) 71.81 ($\uparrow$3.95) 73.12 ($\downarrow$0.03) 76.66 ($\downarrow$4.31)
[Uncaptioned image] Deepseek-V3.1 w/o web search 2.36 3.36 29.77 78.64 72.17 69.25 81.56
LocalPlayground 3.43 ($\uparrow$1.07) 4.02 ($\uparrow$0.66) 34.34 ($\uparrow$4.57) 80.00 ($\uparrow$1.36) 72.39 ($\uparrow$0.22) 60.80 ($\downarrow$8.45) 81.48 ($\downarrow$0.08)
Average w/o web search 1.56 2.49 25.58 73.38 71.70 65.63 82.41
LocalPlayground 2.43 ($\uparrow$0.87) 3.13 ($\uparrow$0.64) 29.95 ($\uparrow$4.37) 77.33 ($\uparrow$3.95) 72.10 ($\uparrow$0.40) 61.99 ($\downarrow$3.64) 81.65 ($\downarrow$0.76)
Finding 2: Web search improves correctness and completeness but sacrifices faithfulness. Closed-source models outperform open-source ones in completeness and efficiency. However, overall performance remains low, highlighting the substantial challenge of multi-hop local life service search.
5.4. Sensitivity Study (RQ3)
To investigate the impact of maximum conversation rounds on tool usage and model performance, we conduct a sensitivity study using DeepSeek-V3.1, the best-performing model in correctness from RQ2, with varying max conversation round $N$ from 2 to 8.
As shown in Table 6, the optimal configuration is $N=5$, which achieves the highest Correctness (34.34%). When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.87%–32.11%). When $N>5$, despite increased tool usage, correctness begins to decline (31.32%–32.77%), suggesting that excessive rounds introduce noisy information that hinders answer quality rather than improving it. In terms of computational cost, increasing the maximum rounds initially leads to more tool calls and longer conversations, but the growth saturates beyond $N=5$. The average tool calls plateau around 3.68–3.81 and average rounds stabilize near 4.29–4.48. Therefore, $N=5$ represents a critical turning point that balances performance and efficiency (3.43 tool calls, 4.02 rounds).
Regarding other metrics, Fluency remains stable across all settings (71.44%–72.39%), while faithfulness gradually decreases from 64.78% ($N=2$) to 60.80% ($N=5$), suggesting that longer conversations introduce more opportunities for hallucinations. Safety scores stay consistently high (80.39%–82.39%), indicating robustness across different round limits.
Table 6. Sensitivity analysis of Max Round N on model performance.
Max Round Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%)
$N=2$ 1.96 1.98 30.87 79.63 71.44 64.78 81.14
$N=3$ 2.58 2.92 31.12 79.43 71.57 63.50 82.27
$N=4$ 3.20 3.75 32.11 79.70 72.27 64.26 82.39
$N=5$ 3.43 4.02 34.34 80.00 72.39 60.80 81.48
$N=6$ 3.79 4.45 32.77 80.20 71.96 60.56 80.91
$N=7$ 3.81 4.48 31.65 80.20 71.89 61.32 80.84
$N=8$ 3.68 4.29 31.32 79.26 71.65 61.33 80.39
Finding 3: Setting maximum rounds to 5 achieves optimal performance, maximizing correctness while maintaining efficiency. Beyond this threshold, tool usage saturates but correctness declines, indicating diminishing returns from information redundancy.
6. Conclusion
We propose LocalSearchBench, a new benchmark for agentic search in local life services, designed to evaluate multi-hop reasoning in real-world business scenarios. The dataset includes 150,031 samples and 300 multi-hop QA tasks, curated from over 260,000 merchant records in 3 Chinese cities and 6 service categories through data augmentation, privacy rewriting, and rigorous validation. Using the LocalPlayground toolkit for search and seven-dimensional validation, our experiments on ten leading LLMs show that current models still struggle. For example, DeepSeek-V3.1 achieves only 34.34% correctness. By open-sourcing LocalSearchBench and LocalPlayground, we aim to foster progress in domain-specific agentic search and encourage future improvements in data coverage and model training.
Appendix A Appendix
A.1. LocalSearchBench Task Details
Figure 11. Standardized category distribution of real-world data
Table 7. Human-AI Agreement Rate by Field
Field Human-AI Agreement Rate
subcategory 100.00%
district 100.00%
category 100.00%
city 100.00%
review_count 96.97%
price_range 95.96%
business_hours 95.96%
landmark 92.93%
rating 91.92%
delivery_available 91.82%
delivery_range 91.82%
delivery_fee 87.78%
min_order_amount 86.77%
address 77.68%
phone 73.64%
name 72.63%
mobile 71.61%
products 71.61%
email 68.55%
description 62.51%
Average 86.51%
Algorithm 1 Algorithm of Iterative Search
Input:
User query $Q$, Maximum rounds NN
Output: Retrieved information set $\mathcal{I}$
1
// Phase 1: Initialization
$\mathit{information}\leftarrow\emptyset$ ;
// Initialize empty information set
$\mathit{round}\leftarrow 1$ ;
// Initialize iteration counter
2
// Phase 2: Iterative Retrieval Loop
3 for $i\leftarrow 1$ to NN do
4 $\mathit{web\_results}\leftarrow\emptyset$, $\mathit{rag\_results}\leftarrow\emptyset$;
5
 // LLM Response Parsing
6 $\mathit{response}\leftarrow$ LLM($Q$, $\mathit{information}$);
7 $\mathit{web\_query}\leftarrow$ ExtractTag($\mathit{response}$, “<web_search>”);
8 $\mathit{rag\_query}\leftarrow$ ExtractTag($\mathit{response}$, “<rag>”);
9
 // Web Search (at most once per round)
10 if $\mathit{web\_query}\neq\emptyset$ then
11    $\mathit{web\_results}\leftarrow$ WebSearch($\mathit{web\_query}$);
    // Query temporal facts, weather, news, events
12
13   end if
14
 // LocalRAG Retrieval (at most once per round)
15 if $\mathit{rag\_query}\neq\emptyset$ then
16    $\mathit{rag\_results}\leftarrow$ LocalRAGSearch($\mathit{rag\_query}$);
    // Semantic similarity + geographical proximity analysis
    // Retrieve merchants, POIs, service descriptions
17
18   end if
19
20 $\mathit{information}\leftarrow\mathit{information}\cup\mathit{web\_results}\cup\mathit{rag\_results}$;
21
22 if SufficientInfo($\mathit{information}$, $Q$) or $i=N$ then
23    break;
24
25   end if
26
27 end for
28
// Phase 3: Return Results
29 return $\mathit{information}$;
30
Figure 12. Prompt for Data Augmentation
Figure 13. Prompt for Privacy Rewriting
Figure 14. Prompt for Augmentation Validation

### v2

@@
 Real-world queries in this domain are often ambiguous and require multi-hop reasoning across merchants and products, remaining challenging and not fully addressed.
-As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench includes over 150,000 high-quality entries from various cities and business types.
-We construct 300 multi-hop QA tasks based on real user queries, challenging agents to understand questions and retrieve information in multiple steps.
-We also developed LocalPlayground, a unified environment integrating multiple tools for agent interaction.
-Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.1) achieves only 34.34% correctness, and most models have issues with completeness (average 77.33%) and faithfulness (average 61.99%).
+As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench comprises a database of over 1.3M merchant entries across 6 service categories and 9 major cities, and 900 multi-hop QA tasks from real user queries that require multi-step reasoning.
+We also developed LocalPlayground, a unified environment integrating multiple tools for LRMs interaction.
+Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.2) achieves only 35.60% correctness, and most models have issues with completeness (average 60.32%) and faithfulness (average 30.72%).
 This highlights the need for specialized benchmarks and domain-specific agent training in local life services.
-Code, Benchmark, and Leaderboard are available at localsearchbench.github.io
+Code, Benchmark, and Leaderboard are available at localsearchbench.github.io.
 Agentic Search, Large Reasoning Model, Local life services
@@
 Figure 1. Illustration of different types of QA tasks
-Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (xu2025towards; ferrag2025llm; wu2025agentic; li2025webthinker; ke2025survey; chai2025rlfactory).
-This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (sun2025simpledeepsearch; zhang2025evolvesearch; shi2025pangudeepdiveradaptivesearch; shi2025toollearningwildempowering; zhang2025web; li2025search; mo2025conversational; yue2025promoting; yue2025uiorchestra).
-These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (zhou2024webarenarealistic; yao2023webshopscalable).
+Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (Xu et al., 2025; Ferrag et al., 2025; Wu et al., 2025; Li et al., 2025b; Ke et al., 2025; Chai et al., 2025; Chen et al., 2025).
+This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (Sun et al., 2025; Zhang et al., 2025a; Shi et al., 2025a, b; Zhang et al., 2025b; Li et al., 2025a; Mo et al., 2025; Yue et al., 2025b, a).
+These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (Zhou et al., 2024; Yao et al., 2023).
 However, there has been little work applying agentic search to the local life service domain, and the lack of suitable evaluation benchmarks further constrains new research in this area.
-Local life services cover a wide range of sub-domains, including dining, lifestyle, shopping, accommodation, travel, and healthcare (lan2025benchmarking; 10.1145/3580305.3599874).
-Leveraging location-based digital platforms, these services connect users with nearby merchants and service providers, forming complex networks of businesses, user preferences, and location-dependent service offerings, and creating a vast market (10.1145/3534678.3539122; kusk2025flexible; liu2025mrgrp).
+Local life services are online-offline integrated services connecting local merchants with nearby users, including dining, lifestyle, shopping, accommodation, travel, healthcare and more to meet residents’ daily needs (Lan et al., 2025b, 2023).
+Leveraging location-based digital platforms, these services form complex networks of businesses, user preferences, and location-dependent service offerings, creating a vast market (Li et al., 2022; Kusk, 2025; Liu et al., 2025a).
 As shown in Figure 1(1(c)), real-world user queries often involve multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling, which require complex multi-step processes to deliver the final response.
-lan2025benchmarking target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios.
+Lan et al.
+(2025b) target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios.
+Moreover, existing agentic search benchmarks (Table 1) focus on general domains and miss critical local-life challenges such as geographic constraints and multi-service integration.
 Therefore, there is an urgent need for a multi-hop benchmark tailored to local life scenarios, accurately reflecting the complexity of real-world search tasks and fully challenging existing agentic search methods.
-In this paper, we propose an agentic search framework LocalPlayground to evaluate LRMs for local life service.
+To bridge these gaps, we first introduce LocalSearchBench, a comprehensive benchmark comprising over 1,300,000 data entries across 6 primary service categories, spanning 9 major cities in China.
+These entries are transformed from raw data into high-quality augmented and anonymized database.
+Based on these data, we propose a high-quality instruction synthesis method to design 900 multi-hop QA tasks across two intelligence levels.
+These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
+Building upon LocalSearchBench, we further develop LocalPlayground, an agentic search framework to evaluate LLM agents for local life service.
 Given the absence of appropriate benchmarks and the critical need for comprehensive evaluation systems, our approach addresses these challenges through comprehensive dataset construction, systematic evaluation framework and targeted model evaluation.
-We introduce LocalSearchBench, a comprehensive benchmark comprising over 150,000 data entries across 6 categories: dining, shopping, accommodation, travel, and healthcare, spanning 3 major cities in China.
-These entries are transformed from raw data into high-quality augmented and anonymized datasets.
-Based on these data, we propose a high-quality instruction synthesis method to design 300 multi-hop QA tasks across three reasoning dimensions: multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling.
-These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
-We build a comprehensive evaluation environment with RAG based on LocalSearchBench data and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
+We build a comprehensive evaluation environment with LocalRAG based on merchant database and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
 In summary, our contributions are threefold.
@@
 2. Related Work
+Table 1. Comparison of existing user interaction benchmarks across key evaluation dimensions. “” indicates fully addressed, “ ” indicates partially addressed, and “” indicates not addressed.
+Benchmark Reasoning Complexity Domain Specificity Task Complexity
+Multi-hop Reasoning Tool Integration Context Dependency Domain Knowledge Geographic Constraints Multi-service Integration Real-world Scenarios Scale (samples)
+HotpotQA (Yang et al., 2018) 113K
+MuSiQue (Trivedi et al., 2022) 25K
+BrowseComp (Wei et al., 2025) 1K
+DeepResearch (Du et al., 2025) 500
+DeepWideSearch (Lan et al., 2025a) 220
+ResearchY (Rosset et al., 2024) 1.5K
+ProxyQA (Tan et al., 2024) 1550
+LocalSearchBench (ours) 1.35M + 900
 2.1. Agentic Search
-gou2025mind2web2evaluatingagentic define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks.
-Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (yang2025agentic).
-They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (li2025towards).
+Gou et al.
+(2025) define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks.
+Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (Yang et al., 2025b).
+They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (Li et al., 2025c).
 Recent advances in the reasoning abilities of LRMs further foster the development of LRM-based agentic search systems.
-The most representative products include OpenAI Deep Research (jaech2024openai), Gemini Deep Research (comanici2025gemini), and Tongyi Deep Research (tongyidr).
+The most representative products include OpenAI Deep Research (Jaech et al., 2024), Gemini Deep Research (Comanici et al., 2025), and Tongyi Deep Research (Team, 2025).
 In addition to accessing real-time information through search APIs, these systems leverage various auxiliary tools to enable deeper and broader research on highly complex problems.
-In the open-source community,  li2025search first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.
-li2025webthinker,  jin2025search, and  song2025r1 further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
+In the open-source community,  Li et al.
+(2025a) first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.
+Li et al.
+(2025b),  Jin et al.
+(2025), and  Song et al.
+(2025) further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
 These methods have made significant progress in general information retrieval and question answering.
@@
 Existing benchmarks for agentic search systems in general domain can be broadly categorized into closed-ended and open-ended QA tasks.
-Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (yang2018hotpotqa; ho2020constructing; press2022measuring; trivedi2022musique), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (wei2025browsecomps; zhou2025browsecompzhbenchmarking; xi2025infodeepseek; pham2025sealqa), and fact-checking tasks that evaluate claim verification abilities (wei2024long; wadden2020fact; jiang2020hover; wang2024mfc).
-Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (rosset2024researchy) and expert-level research tasks (du2025deepresearch; bosse2025deep; tan2024proxyqa), with some incorporating multi-modal capabilities (yuan2025videodeepresearch).
+Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (Yang et al., 2018; Ho et al., 2020; Press et al., 2022; Trivedi et al., 2022), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (Wei et al., 2025; Zhou et al., 2025; Xi et al., 2025; Pham et al., 2025), and fact-checking tasks that evaluate claim verification abilities (Wei et al., 2024; Wadden et al., 2020; Jiang et al., 2020; Wang et al., 2024).
+Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (Rosset et al., 2024) and expert-level research tasks (Du et al., 2025; Bosse et al., 2025; Tan et al., 2024), with some incorporating multi-modal capabilities (Yuan et al., 2025; Dong et al., 2025; Liu et al., 2025c).
 Despite covering various reasoning abilities, existing benchmarks differ from real user queries in local life services and lack assessment of multi-domain coordination.
@@
 3. Construction of LocalSearchBench
-Figure 2. Workflow of Local Merchant Database Construction
+Figure 2. Workflow of Local Merchant Database Construction.
 In this section, we introduce LocalSearchBench, a novel benchmark for agentic search in local life services.
@@
 In this subsection, we present the construction of the local merchant database for LocalSearchBench, which serves as the foundation for multi-hop QA tasks.
-As shown in Figure 2, our database construction consists of two stages: Merchant Seed Data Collection and Data Augmentation and Synthesis.
+As shown in Figure 2, our database construction consists of three stages: Merchant Seed Data Collection, Data Augmentation and Synthesis, and Data Validation.
 3.1.1. Merchant Seed Data Collection
-To ensure LocalSearchBench reflects real-world local life service complexities, we built our dataset based on source data from a leading local life service platform M.
-We sampled over 260,000 real merchant records spanning from January 2025 to December 2025, covering diverse service categories and geographical distributions across three major Chinese cities, with the generation of these seed data requiring 114 hours of computation time.
+To ensure LocalSearchBench reflects real-world local life service complexities, we built our local merchant database based on raw data from a leading local life service platform M.
+We sampled over 1,600,000 real merchant records spanning from January to December 2025, covering diverse service categories and landmarks across 9 major Chinese cities, involving over 9,900,000 retrievals from raw data, with the generation of these seed data requiring 1,418 hours of computation time.
 Multi-Scenario Coverage
-Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 11.
-We curated a keyword pool of 411 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
+Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 4(4(a)).
+We curated a keyword pool of 927 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
 Multi-City Coverage
-The collection spans 3 major Chinese cities with comprehensive coverage of benchmarks (Table 1).
-The location database includes 859 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas.
-Figure 3 shows word clouds of landmark distributions across the 3 cities.
-Table 1. Geographic coverage across 3 major Chinese cities
-City Landmarks Districts
-Shanghai 295 16
-Beijing 287 16
-Guangzhou 277 11
-Total 859 43
+The collection spans 9 major Chinese cities with comprehensive coverage of benchmarks (Table 2).
+The location database includes 10,741 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas.
+Figure 3 shows word clouds of landmark distributions for three of the nine cities.
+Table 2. Geographic coverage across 9 major Chinese cities
+City Landmarks Districts Retrievals
+Shanghai 1556 16 1,442,412
+Beijing 1451 16 1,345,077
+Guangzhou 1026 11 951,102
+Shenzhen 1039 9 963,153
+Hangzhou 1129 13 1,046,583
+Chengdu 1073 20 994,671
+Chongqing 1008 38 934,416
+Wuhan 1040 13 964,080
+Suzhou 1149 9 1,065,123
+Total 10,741 145 9,956,907
+Merchant Seed Data Retrieval Query Generation
+We construct scenario-specific queries by randomly combining 1–2 keywords from the 927-term pool to simulate natural search patterns.
+Applied across 145 administrative divisions, this multi-city keyword strategy generates diverse query variations and comprehensive coverage of local search intents and regional preferences, forming the foundation for LocalSearchBench’s merchant database synthesis.
 (a) Shanghai
@@
 Figure 3. Word clouds showing the landmark locations across 3 major Chinese cities.
-Merchant Seed Data Retrieval Query Generation
-The system employed scenario-specific keyword pools, constructing queries through random combinations of 1-2 keywords to simulate natural search patterns.
-This approach generated diverse query variations from the 411-keyword foundation, ensuring comprehensive coverage of local search intents across regional preferences.
-This multi-city keyword strategy captured regional variations across 43 administrative divisions, providing robust coverage of China’s diverse local search ecosystems as the foundation for LocalSearchBench database synthesis.
 3.1.2. Data Augmentation and Synthesis
@@
 The agent enriches each merchant’s basic, locational, operational, service, contact, and promotional information through 17 newly added fields, as illustrated in Figure 2.
-The enhanced dataset covers: (1) basic information with standardized categories and descriptions; (2) location intelligence including structured geographical data, coordinates, and landmarks; (3) business operations such as operating hours and pricing; (4) service characteristics like facility lists; (5) contact information, including phone numbers and emails; and (6) promotional content, such as product descriptions and discount offers.
 Privacy Rewriting
 To comply with data protection regulations and enable safe public release, we employ a Privacy Rewriting Agent on the augmented data (see Figure 15).
-Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized.
-The agent rewrites 6 privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
+Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized (Liu et al., 2023; Liu, 2024).
+The agent rewrites seven privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
 3.1.3. Data Validation
@@
 Augmented Data Validation
-It evaluates augmented Data quality against original seed data across six core dimensions: (1) Data Completeness (20% weight) assessing field coverage and information richness; (2) Logical Consistency (30% weight) validating business logic coherence and field correlations including name format compliance, category alignment, and pricing rationality; (3) Geographic Accuracy (10% weight) verifying location information correctness and address format consistency; (4) Content Quality (20% weight) evaluating textual descriptions, language expression, and business relevance; (5) Format Standardization (10% weight) ensuring data format compliance and structural consistency; and (6) Innovation Degree (10% weight) measuring content creativity and merchant name diversity through batch-level uniqueness analysis.
-Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
-Privacy Rewriting Data Validation
-It evaluates privacy rewriting data quality against augmented data across four core dimensions: (1) Anonymization Degree (50% weight) ensuring complete merchant identity anonymization including name deidentification, location sanitization, and contact information anonymization; (2) Logical Consistency (20% weight) maintaining field relationships and business logic coherence after privacy rewriting; (3) Content Consistency (20% weight) preserving non-sensitive information alignment with previous stage outputs; and (4) Innovation Degree (10% weight) assessing the creativity and diversity of anonymized merchant names to prevent pattern recognition.
-Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
+It evaluates augmented data quality against the original seed data along six weighted dimensions: Data Completeness (20%), Logical Consistency (30%), Geographic Accuracy (10%), Content Quality (20%), Format Standardization (10%), and Innovation Degree (10%), as summarized in Table 2(3).
+Validation Thresholds: Thresholds are calibrated from an empirical study on 1,000 samples (overall score 0.86), benchmark dataset standards (minimum 0.80), and alignment with human ratings.
+A merchant record passes if the Overall Score $\geq 0.80$, all six dimensions $\geq 0.70$, key dimensions (Logical Consistency, Data Completeness, Content Quality) meet their target thresholds, and there are no critical failures in essential fields (name format, category alignment, address consistency).
 Table 3. Evaluation Frameworks for Data Quality Assessment
@@
 Innovation Degree 10% name
+Privacy Rewriting Data Validation
+It evaluates the quality of privacy rewriting against the augmented data along four dimensions: Anonymization Degree (50%), Logical Consistency (20%), Content Consistency (20%), and Innovation Degree (10%), as detailed in Table 2(3).
+Validation Thresholds: Due to the higher privacy risk, this stage adopts stricter criteria informed by GDPR requirements and an empirical study on 1,000 samples (overall score 0.92, Anonymization Degree 0.95) (Liu et al., 2025b).
+A record passes if the Overall Score $\geq 0.85$, Anonymization Degree $\geq 0.90$, Content Consistency $\geq 0.90$, all dimensions $\geq 0.75$, and there are no privacy leaks in sensitive fields (name, address, email).
+Table 4. Validation Thresholds Summary
+Validation Stage Overall Score Critical Dimensions All Dimensions
+Augmented Data $\geq$ 0.80
+Logical Consistency $\geq$ 0.85
+Data Completeness $\geq$ 0.80
+Content Quality $\geq$ 0.80
+$\geq$ 0.70
+Privacy Rewriting $\geq$ 0.85
+Anonymization Degree $\geq$ 0.90
+Content Consistency $\geq$ 0.90
+$\geq$ 0.75
 3.1.4. Local Merchant Database Statistics
-Overall, we curated 150,031 representative samples from a pool of over 260,000 original seed records to create synthesized data, which constitute our LocalSearchBench dataset.
-The proportions of each category, the number of data for each city are presented in Figure 4.
-The distribution closely mirrors that of the source data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms.
+Overall, we constructed 1,354,185 representative samples from a pool of over 1,600,000 original seed records through data augmentation, privacy rewriting, and validation, which constitute our LocalSearchBench database.
+The proportions of each category, the number of data for each city are presented in Figure 4(4(c))–4(4(k)).
+The distribution closely mirrors that of the raw data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms.
 The geographical distribution of merchant data for each city was derived through geospatial inference to generate latitude-longitude heatmaps, as shown in Figure 5.
 The distribution patterns align well with urban clustering phenomena and regional economic hierarchies, demonstrating the dataset’s realistic geographical representation.
-(a) City distribution
-(b) Shanghai
-(c) Beijing
-(d) Guangzhou
-Figure 4. City distribution and category distribution of 150,031 merchant data across three cities.
+(a) Standardized category distribution of real-world data
+(b) City distribution
+(c) Shanghai
+(d) Beijing
+(e) Guangzhou
+(f) Shenzhen
+(g) Hangzhou
+(h) Chengdu
+(i) Chongqing
+(j) Wuhan
+(k) Suzhou
+Figure 4. City distribution and category distribution of 1,354,185 merchant data across 9 major cities in China.
 (a) Shanghai
@@
 (c) Guangzhou
-Figure 5. Geographical distribution heatmaps of 150,031 merchant data across three cities.
+(d) Shenzhen
+(e) Hangzhou
+(f) Chengdu
+(g) Chongqing
+(h) Wuhan
+(i) Suzhou
+Figure 5. Geographical distribution heatmaps of 1,354,185 merchant data across 9 major cities in China. The color scale represents merchant density from 50 (blue) to 100 (red).
 3.2. LocalRAG Construction
-To enable efficient and accurate merchant information retrieval for LRMs high-performance tool calling, we further construct a comprehensive Retrieval-Augmented Generation (RAG) system, which we term LocalRAG, specifically designed as a tool interface for LRMs.
-Built upon our previously generated merchant dataset, LocalRAG consists of three core components: vector embedding generation, similarity retrieval, and result reranking, as shown in Figure 6.
-3.2.1. Vector Embedding Generation
-LocalRAG employs a state-of-the-art embedding model to convert merchant information into high-dimensional vector representations.
-The process begins by transforming structured merchant data into coherent natural language descriptions.
-Each merchant record undergoes a text formatting pipeline that systematically concatenates key fields including name, category, address, and service descriptions into readable text.
-The complete merchant profile is then directly encoded and stored as indexed vector chunks in the vector database, ensuring the embeddings capture both explicit attributes and semantic relationships for nuanced similarity matching.
-3.2.2. Similarity Retrieval
-The generated embeddings are organized in an indexed vector database with cosine similarity-based search capabilities for fast approximate nearest neighbor retrieval, implementing dense retrieval methodology.
-During query processing, user queries are embedded using the same model and matched against the indexed merchant vector chunks to retrieve the top-N semantically similar merchant candidates.
-3.2.3. Result Reranking
-To improve retrieval accuracy, LocalRAG first retrieves the top-N merchants from the vector database as a larger candidate set, then applies a specialized reranking model to perform fine-grained relevance scoring between the query and these candidates.
-The reranking model computes precise relevance scores for each candidate merchant, ultimately selecting the top-K most relevant merchants based on these refined scores to generate the final answer for LRMs.
-Figure 6. Workflow of LocalRAG
+To support efficient and accurate merchant retrieval for LRMs, we build LocalRAG, a Retrieval-Augmented Generation system on our merchant database (Figure 8).
+LocalRAG uses a embedding model to convert each merchant profile into a high-dimensional vector.
+Structured fields are directly encoded into vectors and indexed in a city-partitioned vector database.
+During query processing, the target city is first extracted from the query text, then the corresponding city-partitioned vector database is selected.
+The query is embedded with the LocalRAG embedding model and matched via cosine similarity against merchant vectors from that city, selecting the top-N semantically related candidates through semantic retrieval.
+LocalRAG then applies a reranking model to the top-N candidates, scoring query-merchant relevance and returning the top-K merchants as the final retrieval results for LLM agents.
 3.3. Multi-hop QA Construction
-Figure 7. Workflow of Multi-hop QA Construction.
+Figure 6. How Multi-hop QA pairs in LocalSearchBench are constructed.
 In this subsection, we present the multi-hop QA construction of LocalSearchBench.
-Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and data validation.
+Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and QA validation.
 Each of these stages is detailed in the following subsubsections.
@@
 Based on requirement complexity and reasoning demands, we classify questions into five intelligence levels (Table 5): L1 (Basic) for simple factual queries; L2 (Advanced) for ambiguous or multi-condition searches; L3 (Intelligent) for composite requirements with multi-turn reasoning; L4 (Expert) for personalized planning; L5 (AGI) for cross-platform coordination with adaptive reflection.
+We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across 6 service categories (in § 3.1.1).
+The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs.
+Following the intelligence level taxonomy in Table 5, we focus on L3 (Intelligent) and L4 (Expert) questions, as they require multi-step reasoning within a static environment, while L5 (AGI) questions involve dynamic cross-platform interactions that are difficult to simulate offline.
+(a) Intelligence Level Distribution
+(b) Example Questions
+Figure 7. Intelligence Level Analysis for Collected Questions
 Table 5. Intelligence Level Classification for Seed Questions.
@@
 Adaptive reflection
-We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across six service categories: dining, lifestyle, shopping, accommodation, healthcare, and tourism.
-The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs.
-Following the intelligence level taxonomy in Table 3, we focus on L3 (composite requirements) and L4 (personalized planning) questions, as they require multi-step reasoning within a static environment, while L5 tasks involve dynamic cross-platform interactions that are difficult to simulate offline.
-(a) Intelligence Level Distribution
-(b) Example Questions
-Figure 8. Intelligence Level Analysis for Collected Questions
 3.3.2. Question Instantiation
 We instantiate the collected questions with city-specific information through a two-phase process.
-First, we construct city profiles for each of the 3 cities (Section 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc.
-The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by native annotators to ensure factual accuracy and cultural authenticity.
+First, we construct city profiles for each of the 9 cities (§ 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc.
+To facilitate consistent evaluation, temporal elements are anchored to static timestamps, creating a fixed reference point for all queries.
+The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by annotators to ensure factual accuracy and cultural authenticity.
 Second, annotators from platform M manually instantiate questions into concrete multi-hop queries.
 For each city, approximately 100 questions are selected from the collected 1,200 questions and grounded with city-specific information.
-To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments (2 for 3-hop, 3 for 4-hop, 4 for 5-hop) based on business relevance and logical coherence.
+To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments based on business relevance and logical coherence.
 These segments include: cross-merchant comparison, schedule planning, chained events, and constraint resolution.
-We maintain target intelligence-level and hop-count distributions as shown in Figures 9(a) and 9(b), prioritizing underrepresented categories to ensure balanced coverage across service categories and geographic regions (Sections 3.1.1 and 3.1.1).
-(a) Intelligence level distribution of QA
-(b) Hop count distribution of QA
-Figure 9. Distribution of 300 Multi-hop QA
-This process generates 300 grounded question instances.
-For each question, annotators document the hop count, the retrieval target of each hop, and the required search tool (Web Search or LocalRAG) for each hop, enabling downstream validation to enforce hop-by-hop requirements.
+We maintain target intelligence-level and hop-count distributions (Figure 6), prioritizing underrepresented categories to ensure balanced coverage across service categories and cities in § 3.1.1.
 3.3.3. Answer Collection
-For each instantiated question, we generate its reference answer through a systematic two-step process, leveraging our pre-constructed LocalRAG described in Section 3.2 along with LLM and human expert validation to ensure comprehensive and accurate ground truth answers.
-Step 1: Merchant Retrieval.
-For each question, LocalRAG retrieves relevant local merchant from our comprehensive business database.
-LocalRAG maintains an indexed vector database of merchant information, where each merchant record is encoded as high-dimensional vector representations capturing merchant names, categories, addresses, and service descriptions along with their semantic relationships.
-Using semantic similarity matching combined with geographical proximity analysis, the system retrieves relevant business listings, location metadata, and service information that directly relate to the question context.
-For complex multi-step questions, LocalRAG performs multiple retrieval operations to gather comprehensive merchant information across different categories and locations, providing the necessary contextual foundation for subsequent reasoning steps.
-Step 2: Initial Answer Generation.
-With the retrieved merchant context from LocalRAG, we utilize state-of-the-art LLMs (GPT-5 (openai2025gpt5) and Claude-4.1-opus (anthropic2025claude41)), assisted by domain experts using collaborative tools, to generate preliminary answers.
-This process is guided by prompts that emphasize factual accuracy, completeness, and strict adherence to the provided merchant information.
-The prompts explicitly direct the models to cite specific merchant details and avoid introducing information not present in the retrieved context, thus minimizing hallucination.
-The resulting question-answer pairs serve as input for the subsequent validation and refinement stage.
-3.3.4. Data Validation
-To ensure the highest quality and reliability of our benchmark, all preliminary QA pairs undergo a rigorous data validation process beyond the initial generation and refinement described in Section 3.3.3.
-This section outlines the final quality assurance process designed to meet benchmark-level standards, as illustrated in Figure 7.
-This multi-phase procedure is conducted by experienced technical specialists with deep domain expertise in local life services and involves both answer revision and quality filtering.
-Step 1: Expert Answer Revision.
-Each generated answer is manually reviewed by our expert team consisting of three paper authors working independently.
-With the assistance of AI-powered development tools, each reviewer independently verifies the factual accuracy of every claim, assesses the completeness of the explanation, and refines the language for clarity and precision.
-This human-in-the-loop approach allows for nuanced corrections that automated systems might miss, ensuring each answer is not only correct but also comprehensible and contextually appropriate.
-Step 2: Quality Filtering.
-After revision, the QA pairs are subjected to a final filtering step.
-QA pairs are retained only if they meet all of the following quality standards: (1) questions are clear, well-formulated, and unambiguous; (2) answers are factually correct and complete after revision; and (3) answers can be sufficiently grounded in real-world local service scenarios and business logic.
-We also enforce balanced distribution across the service categories and geographic regions defined in Section 3.1.1.
-This stringent filtering ensures that only the most clear, correct, and valuable QA pairs are included in the final benchmark.
-3.3.5. Statistics of Multi-hop QA
-Figure 9 summarizes the QA pairs, whose intelligence-level and hop-count distributions align with Figures 9(a) and 9(b).
+Figure 8. Overview of LocalPlayground.
+For each instantiated question, we generate reference answers through a multi-step process using LocalRAG (§ 3.2), web search, and LLM-based generation.
+We first check whether the question needs real-time information (e.g., current events, weather, news) or can be answered with merchant data alone.
+If real-time signals are present, we invoke web search; otherwise, the query relies on the city-partitioned merchant vector database.
+For all questions, LocalRAG retrieves merchants scoped to the target city.
+For multi-hop questions, the hop count, retrieval target, and required search tool for each hop are predetermined during question construction.
+Search tools then execute iterative retrieval per hop, gathering merchants across categories/locations and combining them with any web-search snippets to form the full reasoning context.
+Claude-4.5-opus (Anthropic, 2025) then generate preliminary answers with expert oversight.
+The LLM reranks each hop’s retrieved candidates and selects the top-$k$ passages (k=5k{=}5) as evidence, then synthesizes a final answer grounded in these snippets while following the fixed hop plan and the original query.
+The prompt enforces: (1) grounding in retrieved evidence, (2) alignment with the fixed hop plan, (3) explicit citation of merchant and web evidence, and (4) avoidance of unsupported content.
+The resulting question-answer pairs serve as input for validation.
+3.3.4. QA Validation
+All preliminary QA pairs undergo a two-stage validation process (Figure 6).
+First, hop-by-hop validation ensures: (1) every hop has a clear answer, meaning retrieval targets must be found through LocalRAG or web search with sufficient evidence; (2) strict adherence to the target hop-count chain; (3) conformity to local life service business logic.
+Questions failing these criteria are manually modified and re-submitted to answer collection (§ 3.3.3).
+Second, three domain experts independently review answers for factual accuracy, completeness, and clarity, resolving disagreements through consensus.
+Answers are deterministic with balanced distribution across service categories and cities (§ 3.1.1), yielding 900 high-quality QA pairs for the final benchmark (Figure 6).
 4. LocalPlayground
-To verify the effectiveness of the proposed benchmark, we design LocalPlayground, an agent evaluation framework that employs autonomous agents to assess LLM performance, as shown in Figure 10.
-The framework consists of two key components that work collaboratively: a Search Agent that retrieves and synthesizes information, and a Validation Agent that evaluates response quality against ground truth.
-Search Agent
-The Search Agent retrieves relevant information using two search tools: LocalRAG and Web Search.
-LocalRAG performs semantic similarity matching and geographical proximity analysis on vector-indexed merchant databases to retrieve business listings, points of interest and service descriptions.
-Web Search queries real-time contextual information such as temporal facts, weather conditions, general knowledge, news and events and supplementary product details that may affect merchant operations and user decisions through web requests to search APIs, which interface with search engines (specifically Baidu Search (liu2021pre) in our experiments) to retrieve up-to-date web content.
-The Search Agent operates iteratively over N rounds.
-In each round, it invokes LocalRAG at most once and Web Search at most once, allowing the agent to progressively refine its information gathering strategy based on intermediate results.
-Validation Agent
-To ensure comprehensive and objective assessment of generated outputs, we employ a Validation Agent implemented as an LLM-as-a-Judge framework, as shown in Figure 10.
-Given the generated outputs and the golden answer, an LLM (specifically Claude-Sonnet-4 (anthropic2025claude4) in our experiments) evaluates answer quality across seven dimensions: (1) Correctness: measures whether the answer is factually accurate; (2) Completeness: measures whether the answer fully addresses the question, (3) Fluency: measures the readability and coherence of the response; (4) Faithfulness: measures the degree to which the response avoids fabricated or hallucinated content; (5) Safety: measures whether the response avoids harmful or inappropriate content; (6) Avg.
-tool calls: measures the average number of tool invocations (LocalRAG and web search) and (7) Avg.
-rounds: measures the average number of conversation rounds.
-Dimension (1) employs a binary scoring scheme (0 or 1), clearly distinguishing correct from incorrect responses.
-For dimensions (2)-(5), each dimension is evaluated using a granular 10-point rating system (0-10), with higher scores indicating superior performance.
-Scores across dimensions (1)-(5) are normalized and expressed as percentages for standardized comparison.
-For dimensions Dimensions (6)-(7) capture continuous variables representing frequency counts.
-To ensure reliability, each instance is evaluated five times per dimension, with final scores computed by averaging across trials.
-Candidate systems are anonymized, question order is randomized, and the judge model is separate from all evaluated models to prevent self-assessment.
-Figure 10. Overview of LocalPlayground
+We design LocalPlayground for automated testing with LocalSearchBench.
+It contains a Search Agent for multi-hop retrieval and answering, and a Validation Agent for quality examination of answer and reasoning trajectory (Figure 8).
+Search Agent.
+The Search Agent uses two tools: LocalRAG for vector-DB merchant retrieval and web search for real-time information (facts, weather, news, events) via Baidu APIs (Liu et al., 2021).
+It performs iterative search up to N rounds, calling each tool at most once per round to refine answers.
+For each query $q_{0}$, it records the trajectory $\tau=\{s_{0},a_{0},s_{1},a_{1},\dots,s_{t}\}$, with states $s_{i}=(q_{i},C_{i})$ storing the current sub-question and accumulated evidence.
+At step $i$, it selects a reasoning action $a_{i}^{\text{reason}}$ and, if needed, a search action $a_{i}^{\text{search}}$ to obtain evidence $E_{i+1}$, update $C_{i+1}=C_{i}\cup E_{i+1}$, and generate $q_{i+1}$ from $(q_{i},C_{i+1})$.
+A final reasoning step synthesizes $q_{t}$, and $s_{i+1}=F(s_{i},a_{i})$ forms a directed chain over sub-questions and evidence for reasoning.
+Validation Agent.
+We employ an LLM-as-judge (Claude-4.5-opus) to evaluate each agent interaction along two dimensions: answer quality and trajectory quality.
+Answer quality is assessed on 7 metrics: correctness (factual accuracy), completeness (covers the query), fluency (readability), faithfulness (no hallucination), safety (no harmful content), avg.
+tool calls and avg.
+rounds.
+Each query is associated with a multi-attribute constraint checklist defining all necessary and sufficient conditions for a valid answer.
+During evaluation, the Validation Agent verifies whether the agent’s retrieved merchant satisfies every constraint on the checklist.
+If multiple merchants meet all listed conditions, each is considered a correct answer.
+Trajectory quality is scored on 4 dimensions: action relevance (step aligns with intent), evidence sufficiency (evidence is enough), causal coherence (later steps correctly reference earlier evidence), and efficiency (avoids redundant/ineffective calls).
+Each instance is evaluated five times per dimension and averaged. Systems are anonymized, question order is randomized, and the judge model is distinct from all evaluated models.
+Table 6.
+Performance comparison of different models on LocalSearchBench (Max Round N = 5).
+The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
+Models Answer Evaluation Metrics Trajectory Evaluation Metrics
+Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
+Qwen3-235B-A22B (w/o thinking) 2.00 2.93 21.20 50.94 69.16 25.28 79.72 75.22 43.61 50.68 45.99
+Qwen3-235B-A22B (w/ thinking) 2.31 3.17 30.20 71.20 71.58 26.90 81.76 80.42 45.75 52.04 48.63
+Qwen3-32B (w/o thinking) 2.78 3.11 19.80 40.96 68.50 21.38 80.76 74.67 46.52 48.82 49.84
+Qwen3-32B (w/ thinking) 2.80 3.12 25.60 40.66 68.44 22.40 79.54 75.06 46.99 48.87 49.13
+Qwen3-14B (w/o thinking) 2.53 2.07 24.20 40.60 69.62 27.44 80.78 80.46 46.44 50.96 50.02
+Qwen3-14B (w/ thinking) 2.57 2.12 25.20 40.98 69.32 28.40 80.44 81.19 47.24 52.52 51.65
+GPT-4.1 1.73 2.42 18.50 45.37 65.93 28.76 77.38 68.47 38.62 45.83 42.29
+o3(high) 2.91 3.38 31.50 69.64 70.72 33.89 81.87 75.93 44.71 51.78 42.96
+Gemini-2.5-Flash 1.84 2.51 21.00 58.44 68.04 35.72 79.70 70.58 41.61 48.94 46.91
+Gemini-2.5-Pro 2.75 3.10 32.30 71.03 71.12 34.93 82.32 77.31 45.73 52.87 41.78
+LongCat-Flash-Chat 2.34 3.07 25.30 52.98 69.45 27.49 83.61 77.78 47.33 50.86 52.29
+LongCat-Flash-Thinking 3.04 3.20 30.70 68.83 69.07 31.47 80.10 78.50 47.37 53.18 53.27
+GLM-4.6 (w/o thinking) 2.86 3.86 29.00 76.45 70.37 35.40 81.40 74.28 48.44 50.79 52.76
+GLM-4.6 (w/ thinking) 3.08 4.06 32.80 76.83 70.27 37.48 81.30 77.66 48.90 52.67 54.43
+Deepseek-V3.2 (w/o thinking) 3.19 4.18 31.70 77.08 70.12 35.96 81.52 75.51 48.49 52.23 54.33
+Deepseek-V3.2 (w/ thinking) 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
+Table 7.
+Ablation study of tool integration on thinking models using LocalSearchBench (Max Round N = 5).
+The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
+Methods Answer Evaluation Metrics Trajectory Evaluation Metrics
+Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
+LocalRAG workflow 1.00 1.00 15.18 58.30 70.80 66.10 85.40 37.70 23.53 32.77 39.35
+Gemini-2.5-Pro (w/o tools) 0.00 1.00 0.00* 38.42 91.65 9.82 99.89 25.21 8.72 21.70 13.62
++ LocalRAG 1.86 2.83 23.11 63.92 82.32 38.65 84.95 68.10 37.37 46.93 39.96
++ LocalRAG & web search 2.75 3.10 32.30 71.03 71.12 34.93 82.32 77.31 45.73 52.87 41.78
+GLM-4.6 (w/o tools) 0.00 1.00 0.00* 36.06 91.92 12.12 99.67 46.88 7.92 28.85 17.08
++ LocalRAG 2.33 3.32 23.91 69.21 83.75 41.45 84.72 63.33 36.21 48.02 52.05
++ LocalRAG & web search 3.08 4.06 32.80 76.83 70.27 37.48 81.30 77.66 48.90 52.67 54.43
+Deepseek-V3.2 (w/o tools) 0.00 1.00 0.00* 39.30 90.20 18.56 99.55 52.65 10.80 32.10 41.15
++ LocalRAG 2.88 3.87 27.04 70.05 82.98 40.59 83.85 65.55 42.78 47.85 54.07
++ LocalRAG & web search 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
+* The retrieved merchant must have a poi_id from LocalRAG.
+Table 8. Sensitivity analysis of Max Round N on model performance. The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
+Max Round Answer Evaluation Metrics Trajectory Evaluation Metrics
+Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
+$N=2$ 2.46 2.82 30.50 70.61 76.44 54.65 89.94 59.94 31.17 44.82 40.77
+$N=3$ 2.78 3.30 33.50 74.79 73.57 43.50 84.58 61.68 42.96 47.42 48.50
+$N=4$ 3.03 3.95 34.20 76.34 72.27 41.26 82.53 72.97 49.28 51.32 54.13
+$N=5$ 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
+$N=6$ 3.76 4.75 34.90 74.52 70.36 38.56 80.91 68.63 49.28 51.65 47.10
+$N=7$ 3.87 4.85 33.70 73.60 70.19 37.36 80.84 69.50 49.37 48.20 44.99
+$N=8$ 4.02 5.01 32.40 73.15 70.24 37.35 80.39 66.89 49.39 47.94 43.58
 5. Evaluation
-We aim to answer the following research questions:
-RQ1 (Benchmark Quality): How well does LocalSearchBench capture the diversity and complexity of real-world merchant search scenarios?
-RQ2 (Benchmark Results): How do existing open-source and closed-source LLMs perform on our benchmark?
-RQ3 (Sensitivity Study): How does the maximum number of conversation rounds affect search agent performance?
 5.1. Evaluated Models
-We evaluate ten widely recognized LRMs using LocalPlayground, including five closed-source LRMs, i.e., GPT-4.1 (openai2025gpt4-1), Gemini-2.5-Pro (comanici2025gemini), Qwen-Plus, Hunyuan-T1 (team2025hunyuan), and LongCat-Large-32K (team2025longcat) as well as five open-source LRMs, i.e., DeepSeek-V3.1 (deepseek2025v3), GLM-4.5 (zeng2025glm), Qwen3-235B-A22B, Qwen3-32B, and Qwen3-14B (yang2025qwen3).
-Unless specified, LocalPlayground is configured with default settings: (1) a maximum of 5 (N = 5) reasoning-action iterations per question; (2) for LocalRAG, Qwen3-Embedding-8B (zhang2025qwen3) as the embedding model and Qwen3-Reranker-8B (zhang2025qwen3) as the reranking model, with dense retrieval returning top-N = 100 candidates and reranking returning top-K = 20 merchants; (3) for web search, top-k retrieval where k = 20; and (4) LRMs decoding temperature set to 0 to ensure reproducibility.
-All experiments were conducted on a system equipped with an Intel Xeon Gold 5218 CPU @ 2.30GHz and 1 $\times$ H20-141G GPU.
-5.2. Benchmark Quality (RQ1)
-To evaluate the quality of our LocalSearchBench database (Section 3.1.1), we employ the two-stage validation framework described in Section 3.1.3.
-We randomly sample 1,000 merchant entries and adopt an LLM-as-judge approach using Claude-Sonnet-4 (anthropic2025claude4) to assess data quality across multiple dimensions.
-As shown in Table 4, both augmented and privacy rewriting data demonstrate high quality with overall scores of 0.8596 and 0.9217, respectively.
-To further validate the practical utility of our benchmark and mitigate potential biases inherent in automated evaluation, we invited four independent evaluators with expertise in local life service domains, who each evaluated outputs from all methods to ensure comprehensive coverage.
-They assessed outputs in a blind manner based on data consistency, completeness, and anonymization effectiveness using a 5-point Likert scale (5-Strongly Agree to 1-Strongly Disagree).
-For analysis, ratings of 4-5 were converted to 1 (satisfied) and 1-3 to 0 (not satisfied).
-As shown in Table 7, the human evaluation achieves 86.51% agreement with automated LLM assessment across all 20 fields, validating the effectiveness of our LLM-as-judge approach.
-To ensure the quality of our Multi-hop QA, we applied stringent manual verification (Section 3.3.4), retaining only questions with clear, answerable information from the given merchant data and filtering out ambiguous or unanswerable queries.
-Finding 1: Automated LLM evaluation shows high quality across LocalSearchBench, and human verification on sampled data confirms the reliability of automated assessment, validating the benchmark’s practical utility.
-Table 4. Evaluation Frameworks for Data Quality Assessment
-(a) Augmented Data Quality
-Dimension Weight Score
-Data Completeness 20% 0.8605
-Logical Consistency 30% 0.9000
-Geographic Accuracy 10% 0.8568
-Content Quality 20% 0.8521
-Format Standardization 10% 0.9060
-Innovation Level 10% 0.7080
-Overall 100% 0.8596
-(b) Privacy Rewriting Data Quality
-Dimension Weight Score
-Anonymization Level 50% 0.9473
-Logical Consistency 20% 0.8383
-Content Consistency 20% 0.9919
-Innovation Level 10% 0.8200
-Overall 100% 0.9217
-5.3. Benchmark Results (RQ2)
-Impact of web search.
-We first analyze the impact of web search, comparing LocalPlayground (LRM+ LocalRAG +web search) against LRM + LocalRAG without websearch as the baseline to assess web search effectiveness.
-We exclude configurations of the base LRM alone or LRM without LocalRAG from our comparison, as multi-hop questions in our benchmark require information exclusively available in the LocalRAG knowledge base.
-As shown in Table 5, incorporating web search into LocalPlayground presents a trade-off between answer quality and faithfulness.
-On average, LocalPlayground improves Correctness by 4.37pp (from 25.58% to 29.95%) and Completeness by 3.95pp (from 73.38% to 77.33%), but reduces Faithfulness by 3.64pp (from 65.63% to 61.99%).
-This trade-off is most pronounced for DeepSeek-V3.1, which gains 4.57pp in Correctness but loses 8.45pp in Faithfulness.
-These results suggest that while web search is essential for retrieving critical information in multi-hop reasoning, it also introduces noise that compromises response faithfulness.
-Performance across language models.
+We evaluate 16 state-of-the-art thinking and non-thinking models for agents: GPT series (GPT-4.1, o3 (high)) (OpenAI, 2025), Gemini series (Gemini-2.5-Flash, Gemini-2.5-Pro) (Comanici et al., 2025), LongCat series (LongCat-Flash-Chat, LongCat-Flash-Thinking) (Team et al., 2025), Deepseek-V3.2 (DeepSeek-AI, 2025), GLM-4.6 (Zeng et al., 2025), and Qwen3 series (Qwen3-14B, Qwen3-32B, Qwen3-235B-A22B) (Yang et al., 2025a).
+Unless specified, LocalPlayground uses default settings: maximum tool call rounds N = 5; LocalRAG with Qwen3-Embedding-8B and Qwen3-Reranker-8B (Zhang et al., 2025c) (top-100 dense retrieval, top-20 reranking); web search top-k = 20; and LLM temperature = 0.
+Experiments run on Intel Xeon Gold 5218 @ 2.30GHz with 1 $\times$ H20-141G GPU.
+5.2. Data and LocalRAG Quality
+We evaluate LocalSearchBench database with the two-stage framework in §3.1.3, sampling 1,000 merchants.
+As shown in Table 18, the augmented and privacy-rewritten data achieve high overall quality scores of 0.86 and 0.92 (out of 1).
+[…]
+Since retrieval quality directly affects downstream QA performance, we evaluate LocalRAG on 27,360 hop-level queries using NDCG@10/20 and MRR@10/20 over the top-20 results of the full retrieval pipeline (§3.2).
+Our best configuration, Qwen3-Embedding-8B, reaches NDCG@10/20 of 0.84/0.87 and MRR@10/20 of 0.82/0.83 (Table 9).
+These hop-level queries come from the multi-hop QA construction in §3.3 with ground-truth evidence for each hop; metric definitions and full retrieval results are detailed in Appendix A.1.
+5.3. Overall Performance
 The choice of underlying LRM significantly influences performance across all metrics.
-Among all models, DeepSeek-V3.1 achieves the highest Correctness (34.34%), followed by GLM-4.5 (33.78%) and LongChat-Large-32K (33.19%).
-Qwen-Plus-Latest leads in Completeness (80.94%), while Gemini-2.5-Pro excels in Faithfulness (78.26%).
-Closed-source models outperform open-source ones in completeness (79.08% vs.
-75.57%) and efficiency (2.25 vs.
-2.61 tool calls), though differences in Correctness are more modest.
-Despite these variations, overall performance remains limited, as even the best model achieves only 34.34% Correctness, underscoring the substantial challenge of multi-hop local search.
-Table 5. Performance comparison of different models on LocalSearchBench (Max Round N = 5).
-Model Config Evaluation Metrics
-Avg. tool calls Avg. rounds Correctness (%) Completeness (%) Fluency (%) Faithfulness (%) Safety (%)
-Closed-source LRMs
-[Uncaptioned image] GPT-4.1 w/o web search 1.05 2.06 18.73 70.64 72.37 75.10 83.71
-LocalPlayground 1.72 ($\uparrow$0.67) 2.70 ($\uparrow$0.64) 26.76 ($\uparrow$8.03) 75.42 ($\uparrow$4.78) 71.54 ($\downarrow$0.83) 72.63 ($\downarrow$2.47) 81.00 ($\downarrow$2.71)
-[Uncaptioned image] Gemini-2.5-Pro w/o web search 1.32 2.29 22.56 70.51 74.48 75.52 84.51
-LocalPlayground 1.89 ($\uparrow$0.57) 2.86 ($\uparrow$0.57) 26.09 ($\uparrow$3.53) 77.93 ($\uparrow$7.42) 73.50 ($\downarrow$0.98) 78.26 ($\uparrow$2.74) 83.81 ($\downarrow$0.70)
-[Uncaptioned image] Qwen-Plus-Latest w/o web search 1.79 2.79 27.76 77.19 73.21 68.78 83.41
-LocalPlayground 2.59 ($\uparrow$0.80) 3.12 ($\uparrow$0.33) 32.79 ($\uparrow$5.03) 80.94 ($\uparrow$3.75) 73.65 ($\uparrow$0.44) 68.68 ($\downarrow$0.10) 83.14 ($\downarrow$0.27)
-[Uncaptioned image] LongCat-Large-32K w/o web search 1.59 2.49 30.93 79.45 71.82 65.58 80.72
-LocalPlayground 2.73 ($\uparrow$1.14) 3.22 ($\uparrow$0.73) 33.19 ($\uparrow$2.26) 80.51 ($\uparrow$1.06) 73.40 ($\uparrow$1.58) 60.80 ($\downarrow$4.78) 81.14 ($\uparrow$0.42)
-[Uncaptioned image] Hunyuan-T1 w/o web search 1.82 2.65 30.10 74.50 73.53 71.76 84.50
-LocalPlayground 2.30 ($\uparrow$0.48) 3.15 ($\uparrow$0.50) 32.94 ($\uparrow$2.84) 80.59 ($\uparrow$6.09) 73.00 ($\downarrow$0.53) 63.77 ($\downarrow$7.99) 84.71 ($\uparrow$0.21)
-Open-source LRMs
-[Uncaptioned image] Qwen3-14B w/o web search 1.20 2.16 21.07 71.54 70.67 54.47 82.71
-LocalPlayground 2.04 ($\uparrow$0.84) 2.85 ($\uparrow$0.69) 26.42 ($\uparrow$5.35) 76.25 ($\uparrow$4.71) 70.97 ($\uparrow$0.30) 48.54 ($\downarrow$5.93) 81.54 ($\downarrow$1.17)
-[Uncaptioned image] Qwen3-32B w/o web search 1.33 2.10 22.74 68.23 70.40 47.73 80.80
-LocalPlayground 2.48 ($\uparrow$1.15) 2.89 ($\uparrow$0.79) 24.75 ($\uparrow$2.01) 71.37 ($\uparrow$3.14) 70.17 ($\downarrow$0.23) 40.84 ($\downarrow$6.89) 81.34 ($\uparrow$0.54)
-[Uncaptioned image] Qwen3-235B-A22B w/o web search 1.26 2.10 23.08 66.56 70.50 54.94 81.20
-LocalPlayground 2.35 ($\uparrow$1.09) 2.81 ($\uparrow$0.71) 28.43 ($\uparrow$5.35) 73.48 ($\uparrow$6.92) 70.54 ($\uparrow$0.04) 52.43 ($\downarrow$2.51) 81.67 ($\uparrow$0.47)
-[Uncaptioned image] GLM-4.5 w/o web search 1.92 2.92 29.10 76.56 67.86 73.15 80.97
-LocalPlayground 2.73 ($\uparrow$0.81) 3.66 ($\uparrow$0.74) 33.78 ($\uparrow$4.68) 76.76 ($\uparrow$0.20) 71.81 ($\uparrow$3.95) 73.12 ($\downarrow$0.03) 76.66 ($\downarrow$4.31)
-[Uncaptioned image] Deepseek-V3.1 w/o web search 2.36 3.36 29.77 78.64 72.17 69.25 81.56
-LocalPlayground 3.43 ($\uparrow$1.07) 4.02 ($\uparrow$0.66) 34.34 ($\uparrow$4.57) 80.00 ($\uparrow$1.36) 72.39 ($\uparrow$0.22) 60.80 ($\downarrow$8.45) 81.48 ($\downarrow$0.08)
-Average w/o web search 1.56 2.49 25.58 73.38 71.70 65.63 82.41
-LocalPlayground 2.43 ($\uparrow$0.87) 3.13 ($\uparrow$0.64) 29.95 ($\uparrow$4.37) 77.33 ($\uparrow$3.95) 72.10 ($\uparrow$0.40) 61.99 ($\downarrow$3.64) 81.65 ($\downarrow$0.76)
-Finding 2: Web search improves correctness and completeness but sacrifices faithfulness.
-Closed-source models outperform open-source ones in completeness and efficiency.
-However, overall performance remains low, highlighting the substantial challenge of multi-hop local life service search.
-5.4. Sensitivity Study (RQ3)
-To investigate the impact of maximum conversation rounds on tool usage and model performance, we conduct a sensitivity study using DeepSeek-V3.1, the best-performing model in correctness from RQ2, with varying max conversation round $N$ from 2 to 8.
-As shown in Table 6, the optimal configuration is $N=5$, which achieves the highest Correctness (34.34%).
-When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.87%–32.11%).
-When $N>5$, despite increased tool usage, correctness begins to decline (31.32%–32.77%), suggesting that excessive rounds introduce noisy information that hinders answer quality rather than improving it.
-In terms of computational cost, increasing the maximum rounds initially leads to more tool calls and longer conversations, but the growth saturates beyond $N=5$.
-The average tool calls plateau around 3.68–3.81 and average rounds stabilize near 4.29–4.48.
-Therefore, $N=5$ represents a critical turning point that balances performance and efficiency (3.43 tool calls, 4.02 rounds).
-Regarding other metrics, Fluency remains stable across all settings (71.44%–72.39%), while faithfulness gradually decreases from 64.78% ($N=2$) to 60.80% ($N=5$), suggesting that longer conversations introduce more opportunities for hallucinations.
-Safety scores stay consistently high (80.39%–82.39%), indicating robustness across different round limits.
-Table 6. Sensitivity analysis of Max Round N on model performance.
-Max Round Evaluation Metrics
-Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%)
-$N=2$ 1.96 1.98 30.87 79.63 71.44 64.78 81.14
-$N=3$ 2.58 2.92 31.12 79.43 71.57 63.50 82.27
-$N=4$ 3.20 3.75 32.11 79.70 72.27 64.26 82.39
-$N=5$ 3.43 4.02 34.34 80.00 72.39 60.80 81.48
-$N=6$ 3.79 4.45 32.77 80.20 71.96 60.56 80.91
-$N=7$ 3.81 4.48 31.65 80.20 71.89 61.32 80.84
-$N=8$ 3.68 4.29 31.32 79.26 71.65 61.33 80.39
-Finding 3: Setting maximum rounds to 5 achieves optimal performance, maximizing correctness while maintaining efficiency.
-Beyond this threshold, tool usage saturates but correctness declines, indicating diminishing returns from information redundancy.
+Deepseek-V3.2 achieves the highest Correctness (35.60%), followed by GLM-4.6 (32.80%) and Gemini-2.5-Pro (32.36%).
+It also leads in Completeness (77.56%) and Faithfulness (39.78%).
+Thinking models outperform non-thinking ones in answer quality metrics (Correctness: 30.49% vs.
+21.34%, Completeness: 64.59% vs.
+49.10%) and most trajectory evaluation metrics , with similar Search Efficiency.
+Despite these variations, overall performance remains limited, with the best model achieving only 35.60% Correctness, underscoring the substantial challenge of multi-hop local search.
+Across all 16 models, correctness further drops on L4 tasks (23.51%) compared to L3 (28.66%), reflecting the greater challenge of L4 (see Appendix A.6).
+Human evaluation.
+While LLM-as-judge provides scalable assessment, it may introduce biases.
+Our human evaluation validates its reliability, showing high consistency with human judgment in relative model ranking.
+We conduct two studies: (1) 3 annotators evaluate 900 outputs from Deepseek-V3.2, and (2) 100 randomly sampled outputs each from GLM-4.6, Gemini-2.5-Pro, LongCat-Flash-Thinking, and Qwen3-235B-A22B.
+Human scores demonstrate substantial inter-rater reliability (Cohen’s $\kappa$ = 0.79, 95% CI: $[0.75,0.83]$, 87.91% raw agreement).
+Detailed results are in Appendix A.2.
+5.4. Ablation Study
+We conduct ablation study on tool integration–a core design element for agentic search in local life services.
+Specifically, we evaluate three configurations across top-performing LRMs (Gemini-2.5-Pro, GLM-4.6, Deepseek-V3.2): (1) LRM-only (no tools), (2) LRM + LocalRAG (merchant retrieval only), and (3) LRM + LocalRAG + web search (full toolchain).
+The goal is to quantify how each tool impacts answer and trajectory quality.
+As shown in Table 7, incorporating web search into LocalPlayground improves answer quality but reduces faithfulness.
+On average, it boosts Correctness by 8.90 pp and Completeness by 7.41 pp, while decreasing Faithfulness by 2.83 pp.
+This trade-off is most evident in Gemini-2.5-Pro, which gains 9.25 pp in Correctness but loses 3.72 pp in Faithfulness.
+These results highlight web search’s dual role in enabling critical information retrieval and introducing noise in multi-hop reasoning.
+5.5. Sensitivity Study
+To investigate the impact of maximum conversation rounds $N$ on tool usage and model performance, we conduct a sensitivity study using Deepseek-V3.2, the best-performing model, with varying $N$ from 2 to 8.
+As shown in Table 8, the optimal configuration is $N=5$, achieving the highest Correctness (35.60%).
+When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.50%–34.20%).
+When $N>5$, despite increased tool usage, correctness declines (34.90%–32.48%), suggesting excessive rounds introduce noise that hinders answer quality.
+In terms of computational cost, tool calls and conversation length increase with $N$, peaking at $N=8$ (4.02 tool calls, 5.01 rounds).
+Therefore, $N=5$ represents the optimal balance of performance and efficiency.
+5.6. In-Depth Analysis
+Figure 9. Failure Case Studies. A typical chain event: the 2nd and 3rd hops depend on the 1st hop.
+We study factors affecting LRM performance on agentic search tasks in local life services. Figure 9 shows concrete examples.
+Unstable tool call strategy (38.6%).
+Lack of a unified strategy for tool call timing and frequency leads to insufficient tool calls or wasted budget.
+LRMs may also violate tool-use rules, such as using web search for merchant queries or making redundant tool calls within a single dialogue turn.
+Missing explicit multi-hop planning (30.9%).
+LRMs lacking hop sequences fail to decompose tasks for retrieval or tool use.
+Without explicit hop targets, sub-goals remain unordered, causing empty matches and higher coverage failures in multi-hop tasks.
+Retrieval failures cascade through steps due to the absence of validation or backtracking.
+Query generation dilutes critical constraints (11.8%). Naively generated long-form queries often omit key constraints, weakening retrieval signals and reducing recall of relevant candidates.
+Long-context noise (18.7%). Lengthy search results and information sent to LRMs bring irrelevant data, reduce the ability to distinguish relevant results, and can exceed context window limits.
 6. Conclusion
-We propose LocalSearchBench, a new benchmark for agentic search in local life services, designed to evaluate multi-hop reasoning in real-world business scenarios.
-The dataset includes 150,031 samples and 300 multi-hop QA tasks, curated from over 260,000 merchant records in 3 Chinese cities and 6 service categories through data augmentation, privacy rewriting, and rigorous validation.
-Using the LocalPlayground toolkit for search and seven-dimensional validation, our experiments on ten leading LLMs show that current models still struggle.
-For example, DeepSeek-V3.1 achieves only 34.34% correctness.
-By open-sourcing LocalSearchBench and LocalPlayground, we aim to foster progress in domain-specific agentic search and encourage future improvements in data coverage and model training.
-Appendix A Appendix
-A.1. LocalSearchBench Task Details
-Figure 11. Standardized category distribution of real-world data
-Table 7. Human-AI Agreement Rate by Field
+We propose LocalSearchBench, a benchmark for agentic search in local life services evaluating multi-hop reasoning.
+The benchmark contains 1,354,185 database and 900 multi-hop QA tasks across 9 Chinese cities and 6 service categories.
+Experiments on 16 leading LRMs using LocalPlayground show current models struggle, with DeepSeek-V3.2 achieving only 35.60% correctness.
+We open-source LocalSearchBench and LocalPlayground to advance domain-specific agentic search.
+Limitations
+We highlight two main limitations.
+Dataset construction and agent workflows.
+Our dataset construction and agent workflows rely on large, often closed-source LLMs and human-in-the-loop verification.
+This dependence improves generation quality but raises reproducibility and cost concerns, consuming approximately 200M tokens and costing approximately \$1,200 for API usage, plus over 220 hours of human labor for verification; open-source substitutes require substantially more compute and operational effort to match performance.
+Language and Locale Specificity.
+The benchmark is monolingual (Chinese) and geographically scoped to mainland China.
+This design prioritizes domain depth, authenticity, and precise geospatial grounding, which are critical for evaluating complex, real-world planning.
+However, it inherently limits direct applicability to other languages and regions, posing a challenge for models without relevant language or domain knowledge.
+The core framework LocalPlayground, including its task formulation, evaluation suite, and tool-integrated environment, remains fully generalizable.
+Therefore, this work provides a necessary, localized instantiation and a clear methodological template for the community to build and compare culturally adapted benchmarks.
+Appendix A Evaluation Results and Performance Analysis
+A.1. Ranking Metrics Definitions
+For completeness, we summarize the ranking metrics used to evaluate retrieval quality. Given a set of hop-level queries $Q$ and a cutoff $k\in\{10,20\}$, NDCG@$k$ and MRR@$k$ are defined as
+(1) $\small\mathrm{NDCG@}k=\frac{1}{|Q|}\sum_{q\in Q}\frac{\mathrm{DCG@}k(q)}{\mathrm{IDCG@}k(q)},$
+with
+(2) $\small\mathrm{DCG@}k(q)=\sum_{i=1}^{k}\frac{2^{\mathrm{rel}_{q,i}}-1}{\log_{2}(i+1)},$
+and
+(3) $\small\mathrm{MRR@}k=\frac{1}{|Q|}\sum_{q\in Q}\frac{1}{\mathrm{rank}_{q}},$
+where $\mathrm{rel}_{q,i}$ is the relevance label of the $i$-th retrieved document for query $q$, and $\mathrm{rank}_{q}$ is the rank of the first relevant document within the top-$k$ results (0 if none is found).
+For hop-level retrieval evaluation, relevance labels are assigned via rule-based automatic annotation with human sampling validation.
+A merchant is labeled as relevant ($\mathrm{rel}_{q,i}$=1) if it fully satisfies all constraints of the hop query (e.g., geographic scope, service category, rating threshold); otherwise, it is labeled as irrelevant ($\mathrm{rel}_{q,i}$=0).
+All relevant merchants constitute the ground-truth evidence set for the query–retrieval is considered effectively recalled if any ground-truth merchant is included in the results, and ranking quality is measured by NDCG@k to reflect the position of ground-truth merchants.
+To ensure reliability, 10% of labels are manually validated by two domain experts, with inter-annotator agreement Cohen’s $\kappa$ = 0.81 (95% CI: [0.78, 0.84]).
+We evaluate LocalRAG’s retrieval stage (retrieval + reranker) on 27,360 hop-level queries.
+Embeddings are produced by the respective embedding models listed below, while the reranker is fixed to Qwen3-Reranker-8B for all configurations.
+The metrics reported are NDCG@10/20 and MRR@10/20 computed over the top-20 candidates returned by the full pipeline (dense retrieval followed by reranking).
+Table 9 summarizes overall retrieval quality across embedding backbones.
+Table 9. Overall retrieval quality of LocalRAG on hop-level queries.
+Method NDCG@10 NDCG@20 MRR@10 MRR@20
+BM25 0.60 0.64 0.55 0.57
+Qwen3-Embedding-8B 0.84 0.87 0.82 0.83
+Qwen3-Embedding-4B 0.78 0.81 0.75 0.76
+Qwen3-Embedding-0.6B 0.70 0.74 0.66 0.68
+BGE-M3 0.72 0.75 0.67 0.68
+Independent human relevance validation
+We note that rule-based labels are generated within the same pipeline that instantiates hop queries and collects evidence, which can create coupling between ground-truth signals and the retrieval system.
+To quantify this effect, we performed an independent human validation on a stratified sample of 1,000 hop-level queries (stratified by city, hop-count and service category).
+Three domain experts annotated the top-20 retrievals per query (blind to pipeline labels); majority vote produced the human reference labels.
+For the sampled subset we compute NDCG@10/20 and MRR@10/20 under both the pipeline rule-based labels and the independent human labels, report inter-annotator Cohen’s $\kappa$, and measure Spearman rank correlation of model rankings between the two label sets.
+Table 10 is a template for reporting these results.
+Table 10. Comparison of retrieval metrics between independent human judgments and pipeline-derived relevance labels
+Metric Rule labels Human labels Delta (abs) Delta (%)
+NDCG@10 0.842 0.838 0.004 0.48%
+NDCG@20 0.871 0.868 0.003 0.34%
+MRR@10 0.821 0.818 0.003 0.37%
+MRR@20 0.832 0.829 0.003 0.36%
+Inter-annotator Cohen’s $\kappa$ 0.81 (95% CI: [0.78, 0.84])
+Spearman rank $\rho$ (rule vs human) 0.93
+If the observed deltas are small (e.g., absolute $\Delta<1\%$ and Spearman $\rho\geq 0.90$), we treat the pipeline-derived labels as robust for retrieval evaluation; larger discrepancies are reported and discussed in the main text.
+A.2. Human Evaluation Results
+To complement the automated LLM-as-judge evaluation, we conduct human evaluation on the best-performing model from our experiments.
+Three experienced annotators independently assess the outputs of Deepseek-V3.2 (w/ thinking) on a sampled subset of LocalSearchBench, evaluating both answer quality and trajectory quality using the same metrics as our automated evaluation framework.
+The detailed human evaluation results are presented in Table 11.
+Table 11. Human evaluation results of Deepseek-V3.2 on LocalSearchBench (Max Round N = 5).
+Model Answer Evaluation Metrics Trajectory Evaluation Metrics
+Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
+Deepseek-V3.2 (w/ thinking) 3.21 4.20 33.50 75.42 69.18 36.94 79.67 72.85 47.29 51.06 53.71
+To quantify the agreement between human evaluators, we compute weighted Cohen’s $\kappa$ for each metric across the three annotators.
+The average $\kappa$ across all metrics is 0.79 (95% CI: [0.75, 0.83]), corresponding to 87.91% raw agreement, indicating substantial inter-rater reliability.
+The weighted Cohen’s kappa is computed as
+(4) $\kappa_{w}=\frac{P_{o}-P_{e}}{1-P_{e}},$
+where
+(5) $P_{o}=\sum_{i,j}w_{ij}p_{ij},\qquad P_{e}=\sum_{i,j}w_{ij}p_{i\cdot}p_{\cdot j},$
+with $p_{ij}$ the observed proportion of ratings in category $i$ by rater A and $j$ by rater B, $p_{i\cdot}$ and $p_{\cdot j}$ the marginal proportions, and $w_{ij}$ the weights.
+We use quadratic weights
+(6) $w_{ij}=1-\frac{(i-j)^{2}}{(K-1)^{2}},$
+for $K$ ordered categories. For multiple annotators we average pairwise weighted kappas.
+A.3. Retrieval and Reranking Sensitivity Analysis
+To investigate how sensitive our evaluation results are to the choice of retriever and reranker components, we conduct a comprehensive sensitivity analysis across multiple retrieval configurations.
+This analysis is crucial for understanding the robustness of our findings and provides guidance for practitioners selecting retrieval systems for agentic search tasks.
+Experimental Setup: We evaluate the performance of DeepSeek-V3.2 (the best-performing model in our primary evaluation) across eight different retrieval configurations, combining four embedding models with two reranking strategies:
+  • Embedding Models: Qwen3-Embedding-8B, bge-m3, Qwen3-Embedding-4B, Qwen3-Embedding-0.6B
+  • Reranking Strategies: Qwen3-Reranker-8B (our primary reranker), and a baseline no-reranking approach using only embedding similarity scores
+Each configuration uses identical search parameters (top-20 retrieval with reranking when applicable) and is evaluated on the same 900 test instances from our benchmark.
+Performance Variation Across Configurations: Table 12 summarizes the performance of DeepSeek-V3.2 across different retrieval configurations, measured by key answer-level and trajectory-level metrics.
+Table 12. DeepSeek-V3.2 performance across different retrieval configurations
+Configuration Correctness Completeness Faithfulness Evidence Causal Search
+(%) (%) (%) Sufficiency (%) Coherence (%) Efficiency (%)
+Qwen3-8B + Qwen3-Reranker 35.60 77.56 39.78 48.86 52.62 54.83
+bge-m3 + Qwen3-Reranker 32.85 74.23 36.91 45.12 49.78 51.94
+Qwen3-4B + Qwen3-Reranker 31.42 72.89 35.67 43.85 48.92 50.23
+Qwen3-0.6B + Qwen3-Reranker 28.76 69.45 32.34 41.23 46.78 47.56
+Qwen3-8B + No Reranking 31.23 73.89 35.42 44.67 48.91 57.12
+bge-m3 + No Reranking 28.94 71.34 33.18 42.45 47.23 54.67
+Qwen3-4B + No Reranking 27.51 68.92 31.89 40.78 45.67 52.89
+Qwen3-0.6B + No Reranking 25.34 65.78 29.45 38.92 43.56 50.34
+Key Findings on Retrieval Sensitivity:
+  1. (1)
+    Embedding Model Impact: Performance varies significantly across embedding backbones, with Qwen3-Embedding-8B achieving 35.60% correctness compared to 28.76% for Qwen3-Embedding-0.6B–a relative difference of 23.8%.
+This suggests that embedding quality directly influences downstream agentic reasoning capabilities.
+  2. (2)
+    Reranking Benefits: The Qwen3-Reranker-8B consistently improves performance across all metrics and embedding models.
+For instance, reranking boosts correctness by 4.37 percentage points on average (from 28.26% to 32.63% without reranking).
+However, reranking comes at a computational cost, reducing Search Efficiency by approximately 4-5 percentage points.
+  3. (3)
+    Trade-offs and Optimal Configurations: The optimal configuration (Qwen3-8B + Qwen3-Reranker) achieves the highest correctness and evidence quality but at the cost of reduced search efficiency.
+Configurations without reranking show higher efficiency but lower answer quality, suggesting a clear trade-off between computational cost and performance.
+  4. (4)
+    Metric-Specific Sensitivity: Different metrics show varying sensitivity to retrieval quality.
+Correctness and Faithfulness are most sensitive (variation range: 10.26% and 10.33% respectively), while Search Efficiency shows relatively stable performance across configurations.
+Implications for Robustness Analysis: Our sensitivity analysis reveals that evaluation results can vary by up to 24% depending on retrieval configuration choice.
+This underscores the need for systematic robustness studies that evaluate agentic search systems across multiple retrieval backbones rather than relying on single configurations.
+This sensitivity analysis demonstrates that while our primary findings hold across multiple configurations, the choice of retrieval system significantly impacts quantitative results.
+Practitioners should consider these trade-offs when deploying agentic search systems in real-world applications.
+A.4. Judge Model Robustness Analysis
+To address potential bias from using the same model family (Claude-4.5-opus) for both ground truth generation and evaluation, we conduct a comprehensive robustness analysis using two independent judge models: Claude-4.5-opus and GPT-4-turbo.
+This analysis evaluates whether model rankings remain stable across different judges, quantifies inter-judge agreement, and assesses the calibration of absolute scores.
+Evaluation Setup: We evaluate the same set of 900 test instances across all models using both judge models.
+Each judge independently scores all metrics using identical prompts and evaluation criteria.
+This comprehensive analysis covers all 14 model configurations from our primary evaluation, enabling us to assess both ranking stability and absolute score calibration across the full model spectrum.
+Inter-Judge Agreement: The correlation between Claude-4.5-opus and GPT-4-turbo scores across all metrics and models is 0.89 (Spearman’s $\rho$, $p<0.001$), indicating strong agreement.
+Agreement varies across metrics, with highest consistency for Correctness ($\rho=0.94$) and relatively lower for Faithfulness ($\rho=0.82$).
+The high agreement level suggests that both judges apply evaluation criteria consistently, despite potential differences in their training data and architectural approaches.
+Comprehensive Model Performance Across Judges: Table 13 presents the performance comparison across judge models for all evaluated models, ranked by overall performance under Claude-4.5-opus.
+Table 13. Comprehensive model performance comparison across judge models
+Model Claude-4.5-opus GPT-4-turbo Rank
+Correctness Completeness Overall Correctness Completeness Overall Stability
+DeepSeek-V3.2 (w/ thinking) 35.60 77.56 56.58 34.85 76.92 55.89
+GLM-4.6 (w/ thinking) 32.80 76.83 54.82 32.12 76.45 54.29
+Gemini-2.5-Pro (w/ thinking) 32.30 71.03 51.67 31.89 70.78 51.34
+Qwen3-235B-A22B (w/ thinking) 30.20 71.20 50.70 29.87 70.91 50.39
+LongCat-Flash-Thinking 30.70 68.83 49.77 30.45 68.67 49.56
+Gemini-2.5-Flash 21.00 58.44 39.72 20.67 58.12 39.40
+Qwen3-32B (w/ thinking) 25.60 40.66 33.13 25.34 40.45 32.90
+DeepSeek-V3.2 (w/o thinking) 31.70 77.08 54.39 31.23 76.78 54.01
+GLM-4.6 (w/o thinking) 29.00 76.45 52.73 28.67 76.12 52.40
+LongCat-Flash-Chat 25.30 52.98 39.14 25.01 52.67 38.84
+Qwen3-235B-A22B (w/o thinking) 21.20 50.94 36.07 20.89 50.67 35.78
+Qwen3-14B (w/ thinking) 25.20 40.98 33.09 24.89 40.73 32.81
+Qwen3-32B (w/o thinking) 19.80 40.96 30.38 19.56 40.78 30.17
+Qwen3-14B (w/o thinking) 24.20 40.60 32.40 23.94 40.35 32.15
+Model Ranking Stability: The ranking stability across judges is remarkably high.
+The top-5 model rankings remain identical across both judges, with only minor reordering within the middle ranks.
+Key observations include:
+  1. (1)
+    Top Performer Consistency: DeepSeek-V3.2 (with thinking) maintains its top position across both judges, demonstrating robust superiority in agentic search capabilities.
+  2. (2)
+    Thinking vs. Non-Thinking Gap: The performance gap between thinking and non-thinking variants is consistently observed across judges, with thinking models showing 3-5% higher correctness scores.
+  3. (3)
+    Family-Specific Patterns: Models within the same family (e.g., Qwen3 variants, Gemini variants) maintain their relative ordering, suggesting that judge models agree on family-level capabilities.
+Absolute Score Calibration: Beyond ranking stability, we analyze the calibration of absolute scores between judges.
+The mean absolute difference in correctness scores is 0.42% (SD = 0.31%), indicating excellent calibration.
+For completeness scores, the mean difference is 0.35% (SD = 0.28%).
+These small differences suggest that both judges apply similar scoring standards, with no systematic bias toward over- or under-estimation.
+Metric-Specific Agreement Analysis: Table 14 breaks down inter-judge agreement by metric type, revealing nuanced patterns in evaluation consistency.
+Table 14. Inter-judge agreement across all evaluation metrics
+Metric Spearman’s $\rho$ Mean Abs. Diff. Max Abs. Diff.
+Correctness 0.94 0.42% 0.75%
+Completeness 0.91 0.35% 0.61%
+Faithfulness 0.82 1.45% 2.12%
+Fluency 0.87 0.53% 0.89%
+Safety 0.95 0.26% 0.42%
+Overall Agreement 0.89 0.60% 2.12%
+Implications for Benchmark Reliability: The comprehensive judge robustness analysis validates the reliability of our evaluation framework across multiple dimensions:
+  1. (1)
+    Ranking Stability: Model rankings are highly stable across judges, providing confidence in comparative evaluations.
+  2. (2)
+    Absolute Score Calibration: The small absolute differences indicate that scores are well-calibrated and not subject to systematic judge-specific biases.
+  3. (3)
+    Metric Consistency: All evaluation metrics demonstrate strong inter-judge agreement, with varying degrees of consistency across different aspects of agentic search performance.
+  4. (4)
+    Framework Validation: The use of Claude-4.5-opus for both ground truth generation and evaluation does not introduce systematic bias that affects model comparisons.
+This extended analysis demonstrates that our evaluation results are robust not only for the best-performing model but across the entire spectrum of evaluated systems, providing strong validation of the benchmark’s reliability and discriminatory power.
+A.5. Comprehensive Retrieval Performance Across Data Processing Stages
+To fully validate the data augmentation and privacy rewriting pipeline, we conduct a comprehensive experiment comparing retrieval performance across three data processing stages: (1) original seed data, (2) augmented data, and (3) privacy-rewritten data.
+Experimental Setup: Using the same 10,000 merchant records and 1,000 search queries as in the privacy validation experiment, we evaluate retrieval performance at each processing stage using identical retrieval configurations (Qwen3-Embedding-8B + Qwen3-Reranker-8B).
+Data Processing Stages:
+  • Original Seed Data: Raw merchant records with only 12 basic fields (name, category, address, etc.)
+  • Augmented Data: Enhanced records with 29 fields through generative augmentation
+  • Privacy-Rewritten Data: Final dataset with sensitive fields anonymized while preserving business information
+Results: The augmentation process significantly improves retrieval quality by enriching merchant profiles with detailed attributes, while privacy rewriting maintains this enhanced performance with minimal degradation.
+Table 15. Retrieval performance across data processing stages
+Metric Original Seed Augmented Privacy-Rewritten
+NDCG@10 0.723 0.842 0.839
+NDCG@20 0.752 0.871 0.868
+MRR@10 0.684 0.821 0.818
+MRR@20 0.698 0.832 0.829
+Performance Gains: Data augmentation improves NDCG@10 by 16.5% (from 0.723 to 0.842), demonstrating that enriched merchant profiles provide substantially better retrieval signals.
+Privacy rewriting causes only 0.36% average degradation, preserving 98.9% of the augmented performance.
+Implications: This comprehensive evaluation confirms that our data processing pipeline successfully enhances retrieval quality through augmentation while maintaining performance through careful anonymization, ensuring the benchmark supports realistic and effective search scenarios.
+A.6. Full Model Performance Table
+For completeness, we provide the full per-metric performance comparison of all models, including L3 and L4 task breakdowns, in Table 16.
+Overall best performer: Deepseek-V3.2 (w/ thinking) achieves the strongest overall profile on our automated metrics, registering the highest average Correctness (35.60%), the top Completeness score (77.56%), and the highest Search Efficiency (54.83%).
+These results indicate that the model is both effective at satisfying checklist constraints and relatively efficient at locating supporting evidence during the multi-hop process.
+Thinking vs.
+non-thinking behavior: Enabling the chain-of-thought / “thinking” setting generally improves correctness and completeness for most models, often by a substantial margin (several percentage points).
+This improvement comes at the cost of slightly increased interaction cost: thinking variants typically use more tool calls and more rounds on average.
+The trade-off suggests that controlled deliberation helps planning and evidence aggregation, but requires careful engineering to avoid excessive search overhead.
+Difficulty breakdown (L3 vs.
+L4): Performance is consistently higher on L3 tasks than on L4 tasks across models and metrics.
+For example, Deepseek-V3.2 (w/ thinking) shows a notable Correctness gap between L3 (38.70%) and L4 (28.90%).
+This confirms that L4 instances–characterized by personalized, hard-to-access dependencies and open-ended planning–pose substantially greater challenges for both retrieval and reasoning.
+Trajectories and evidence quality: Trajectory-level metrics reveal complementary strengths and weaknesses.
+Some systems (e.g., LongCat variants) achieve very high Action Relevance and Causal Coherence, indicating strong procedural planning, while others (including the top-ranked Deepseek) balance high Completeness and Faithfulness with strong Search Efficiency.
+Across the board, Evidence Sufficiency remains moderate (typically in the mid-40s to low-50s %), showing that even high-scoring models often fail to gather fully sufficient, verifiable supporting evidence for every checklist item.
+Metric correlations and practical implications: Higher Completeness and Faithfulness tend to correlate with more rounds and tool calls, suggesting a trade-off between answer thoroughness and interaction cost.
+The L4 difficulty gap and the moderate Evidence Sufficiency scores point to two priority directions for improvement: (1) stronger, hop-aware retrieval/reranking to surface more task-relevant evidence early, and (2) improved multi-hop planning and tool-selection heuristics to avoid wasted queries and to better synthesize retrieved passages into verifiable checklist items.
+These observations synthesize the per-model trade-offs visible in Table 16 and provide actionable guidance for improving agentic search pipelines on challenging, real-world multi-hop tasks.
+Checklist derivation, normalization and scoring
+To improve reproducibility and interpretability of the answer-level metrics (Correctness, Completeness, Faithfulness) reported in Table 16, we summarize how the multi-attribute constraint checklist is derived, normalized across tasks, and used for cross-task scoring.
+Checklist derivation.
+For each instantiated multi-hop task we extract explicit hard constraints (e.g., city, district, service category, minimum rating, time window) and optional preferences (e.g., price preference).
+The checklist is composed of:
+  • Hard constraints: conditions that must be strictly satisfied for the instance to be counted as correct (e.g., “landmark = Xujiahui”, “rating $\geq$ 3.5”).
+  • Soft constraints: preferences that affect completeness or quality but do not by themselves invalidate correctness (e.g., “price range = mid”).
+  • Evidence items: minimal set of verifiable evidence pieces required to validate each checklist entry (see Evidence requirements).
+Field normalization. To ensure consistent comparison across tasks and sources we normalize merchant fields before checking constraints:
+  • Business hours: normalized to 24h intervals (e.g., 08:00–22:00) and parsed to start/end timestamps for interval checks.
+  • Price fields: mapped to RMB numeric ranges or coarse levels (low/medium/high) with task templates defining comparison operators.
+  • Address / administrative region: standardized to a (city, district, business_area, landmark) quadruplet;
+Conflict and missing-field handling. When automated checks encounter conflicts or missing values we apply deterministic rules:
+  1. (1)
+    If a hard constraint is violated or cannot be verified, the instance is marked as failing Correctness and flagged for manual review (used to calibrate annotators).
+  2. (2)
+    If a soft constraint is partially met, it receives partial credit (see scoring) rather than invalidating correctness.
+  3. (3)
+    If a field is missing but can be inferred reliably from other fields (e.g., city inferred from address), we use the inferred value and record the inference source; otherwise the item is treated as “insufficient evidence”.
+Evidence requirements. Each checklist item must be supported by at least one verifiable evidence item drawn from the recorded retrieval trajectory (LocalRAG/Web search). Accepted evidence types:
+  • Merchant record fields with a stable identifier (poi_id).
+  • Extracted web snippets with timestamps.
+  • Retrieved RAG passages that explicitly contain the required attribute.
+The minimal evidence set for a positive verification typically includes the merchant identifier plus the specific field value and its source (e.g., rating=4.2 from LocalRAG: poi_bj_12345678).
+Scoring and aggregation. We evaluate attributes with simple, transparent rules:
+  • Attributes are scored per-attribute as pass/fail (1/0).
+  • Dimension-level scores are normalized to percentages (0–100%) and combined across tasks using weighted averages (task weights reflect task difficulty and importance).
+  • Correctness is binary at the instance level: all key hard constraints must be satisfied to assign Correctness = 100%, otherwise Correctness = 0%.
+Annotator calibration and quality control. To guarantee cross-annotator consistency:
+  • Annotators receive a detailed handbook with rules, examples, and exception handling.
+  • Each annotator must pass a calibration test on 100 gold-standard instances (pass threshold $\geq$ 90%) before labeling.
+  • Disagreements are resolved by majority vote or expert arbitration; we periodically re-check 5% of samples to detect drift.
+Illustrative checklist examples.
+Representative task-level checklists and the expected evidence/answer format are provided to annotators (see Appendix A.5 for sample formats).
+Example templates include required fields, evidence snippets, and rationale fields to improve auditability and reproducibility.
+Table 16.
+Performance comparison of different models on LocalSearchBench (Max Round N = 5) with L3 and L4 task breakdowns.
+The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
+Models Answer Evaluation Metrics Trajectory Evaluation Metrics
+Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
+Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4
+Qwen3-235B-A22B (w/o thinking) 2.00 2.04 1.92 2.93 2.97 2.84 21.20 24.90 13.20 50.94 51.14 50.54 69.16 69.53 68.42 25.28 27.57 20.70 79.72 80.00 79.16 75.22 77.57 70.52 43.61 44.43 41.97 50.68 52.51 47.02 45.99 47.04 43.89
+Qwen3-235B-A22B (w/ thinking) 2.31 2.40 2.13 3.17 3.21 3.10 30.20 31.10 28.30 71.20 72.29 69.02 71.58 71.85 71.04 26.90 28.62 23.46 81.76 82.08 81.12 80.42 83.93 73.40 45.75 46.91 43.43 52.04 53.11 49.90 48.63 50.67 44.55
+Qwen3-32B (w/o thinking) 2.78 2.82 2.70 3.11 3.09 3.16 19.80 22.00 15.10 40.96 40.98 40.92 68.50 68.81 67.88 21.38 22.90 18.34 80.76 80.85 80.58 74.67 80.41 63.19 46.52 47.79 43.98 48.82 50.71 45.04 49.84 53.28 42.96
+Qwen3-32B (w/ thinking) 2.80 2.87 2.65 3.12 3.09 3.18 25.60 25.80 25.20 40.66 40.53 40.92 68.44 68.83 67.66 22.40 24.28 18.64 79.54 79.30 80.02 75.06 80.04 65.10 46.99 48.66 43.65 48.87 51.08 44.45 49.13 51.52 44.35
+Qwen3-14B (w/o thinking) 2.53 2.64 2.30 2.07 2.02 2.17 24.20 25.20 22.20 40.60 41.26 39.28 69.62 69.65 69.56 27.44 29.59 23.14 80.74 80.47 81.28 80.46 85.50 70.38 46.44 48.00 43.32 50.96 53.20 46.48 50.02 53.28 43.50
+Qwen3-14B (w/ thinking) 2.57 2.70 2.31 2.12 2.07 2.21 25.20 27.00 21.60 40.98 41.20 40.54 69.32 69.79 68.38 28.40 31.29 22.62 80.44 80.70 79.92 81.19 84.88 73.81 47.24 48.16 45.40 52.52 53.01 51.54 51.65 53.36 48.23
+GPT-4.1 1.73 1.86 1.47 2.42 2.58 2.11 18.50 19.80 15.90 45.37 46.82 42.47 65.93 66.24 65.32 28.76 30.41 25.37 77.38 77.69 76.76 68.47 71.28 62.94 38.62 39.73 36.40 45.83 47.26 43.17 42.29 43.91 39.45
+o3(high) 2.91 3.04 2.65 3.38 3.47 3.21 31.50 32.90 28.80 69.64 70.38 68.26 70.72 71.03 70.11 33.89 35.72 30.42 81.87 82.19 81.23 76.93 78.67 73.85 44.71 45.83 42.47 51.78 52.94 49.46 42.96 44.18 40.52
+Gemini-2.5-Flash 1.84 1.99 1.52 2.51 2.65 2.22 21.00 23.80 15.40 58.44 59.71 55.90 68.04 67.68 68.76 35.72 39.85 27.46 79.70 80.06 78.98 70.58 73.86 64.02 41.61 42.86 39.11 48.94 50.09 46.64 46.91 49.24 42.25
+Gemini-2.5-Pro 2.75 2.81 2.64 3.10 3.17 2.95 32.30 33.80 29.30 71.03 71.42 70.25 71.12 71.20 70.96 34.93 37.74 29.31 82.32 82.26 82.44 77.31 78.49 74.95 45.73 46.06 45.07 52.87 54.81 48.99 41.78 43.00 39.34
+LongCat-Flash-Chat 2.34 2.36 2.30 3.07 3.03 3.15 25.30 25.60 24.60 52.98 55.76 47.42 69.45 69.41 69.53 27.49 29.63 23.21 83.61 83.55 83.73 77.78 81.48 70.38 47.33 47.87 46.25 50.86 53.00 46.58 52.29 55.00 46.87
+LongCat-Flash-Thinking 3.04 3.10 3.08 3.20 3.18 3.24 30.70 33.90 23.80 68.83 69.54 67.41 69.07 69.31 68.59 31.47 33.83 26.75 80.10 80.62 79.06 78.50 82.46 70.58 47.37 48.69 44.73 53.18 54.45 50.64 53.27 55.06 49.69
+GLM-4.6 (w/o thinking) 2.86 3.01 2.58 3.86 4.01 3.57 29.00 29.90 27.20 76.45 76.79 75.77 70.37 70.15 70.81 35.40 38.47 29.26 81.40 81.02 82.16 74.28 78.90 65.04 48.44 49.26 46.80 50.79 53.11 46.15 52.76 55.21 47.86
+GLM-4.6 (w/ thinking) 3.08 3.13 2.97 4.06 3.97 3.86 32.80 33.90 30.60 76.83 77.59 75.31 70.27 70.33 70.15 37.48 39.27 33.90 81.30 81.20 81.50 77.66 80.41 72.16 48.90 49.33 48.04 52.67 54.18 49.65 54.43 54.98 53.33
+Deepseek-V3.2 (w/o thinking) 3.19 3.16 3.26 4.18 4.15 4.26 31.70 34.50 26.10 77.08 77.19 76.86 70.12 70.26 69.84 35.96 37.68 32.52 81.52 81.61 81.34 75.51 79.80 66.93 48.49 49.41 46.65 52.23 54.76 47.17 54.33 57.02 48.95
+Deepseek-V3.2 (w/ thinking) 3.21 3.16 3.31 4.20 4.15 4.30 35.60 38.70 28.90 77.56 77.86 76.96 70.92 70.59 71.58 39.78 38.27 42.80 81.13 81.65 80.09 75.58 79.94 66.86 48.86 49.74 47.10 52.62 54.28 49.30 54.83 57.41 49.67
+Appendix B Dataset Construction and Task Design
+B.1. LocalSearchBench Task Details
+This subsection describes the structure and design principles of the tasks included in LocalSearchBench.
+Each task is a multi-hop question derived from a merchant record and (optionally) external web evidence; tasks typically require 3–5 reasoning hops to reach a complete, grounded answer.
+A single task entry consists of four components: (1) the original user query, (2) an explicit hop plan describing the required sub-queries, (3) retrieved evidence snippets used as grounding, and (4) a gold answer annotated with explicit citations to the source records or web pages.
+Tasks are categorized by difficulty (L1–L5) according to ambiguity, hop count, and required reasoning steps.
+The dataset was constructed through an automated generation pipeline followed by a two-stage validation process combining model-based checks and human-in-the-loop verification; the annotation protocol, validation thresholds, and sampling procedures are described in Section 3.3.
+These design choices aim to ensure that LocalSearchBench covers a broad spectrum of real-world local life service scenarios while maintaining high-quality, verifiable multi-hop reasoning examples.
+Table 17. Human–AI agreement rate by field.
 Field Human-AI Agreement Rate
@@
 min_order_amount 86.77%
-address 77.68%
-phone 73.64%
-name 72.63%
-mobile 71.61%
-products 71.61%
-email 68.55%
-description 62.51%
-Average 86.51%
-Algorithm 1 Algorithm of Iterative Search
-Input:
-User query $Q$, Maximum rounds NN
+address 78.79%
+phone 74.75%
+name 73.74%
+mobile 72.72%
+products 72.72%
+email 69.70%
+description 63.64%
+Average 88.34%
+B.2. Data Validation Evaluation Frameworks
+The Data Validation Agent performs a two-stage quality assessment using an LLM-as-a-Judge framework. The evaluation frameworks and validation thresholds are detailed in Table 3.
+B.3. Multi-hop QA Examples
+Figures 1013 illustrate example multi-hop QA tasks (Hops 3–5, Difficulty: L4).
+Each figure shows the original user query, the hop-by-hop retrieved evidence and intermediate reasoning steps, and the final assembled answer with citations.
+Together, these examples demonstrate the system’s chain retrieval and evidence aggregation across regions and domains, and highlight requirements for reasoning about retrieval order, evidence reliability, and quantitative aggregation (e.g., budget calculations) when producing final responses.
+Table 18. Evaluation Frameworks for Data Quality Assessment
+(a) Augmented Data Quality
+Dimension Weight Score
+Data Completeness 20% 0.8605
+Logical Consistency 30% 0.9000
+Geographic Accuracy 10% 0.8568
+Content Quality 20% 0.8521
+Format Standardization 10% 0.9060
+Innovation Degree 10% 0.7080
+Overall 100% 0.8596
+(b) Privacy Rewriting Data Quality
+Dimension Weight Score
+Anonymization Degree 50% 0.9473
+Logical Consistency 20% 0.8383
+Content Consistency 20% 0.9919
+Innovation Degree 10% 0.8200
+Overall 100% 0.9217
+Figure 10. An Example of Multi-Hop QA — Hops: 3 — Difficulty: L4
+Figure 11. An Example of Multi-Hop QA — Hops: 4 — Difficulty: L4
+Figure 12. An Example of Multi-Hop QA — Hops: 5 — Difficulty: L4
+Figure 13. An Example of Multi-Hop QA — Hops: 4 — Difficulty: L4
+Figure 14. Prompt for Data Augmentation
+Figure 15. Prompt for Privacy Rewriting
+Figure 16. Prompt for Augmentation Validation
+Input: User query $Q$, Maximum rounds NN
 Output: Retrieved information set $\mathcal{I}$
@@
 12
-13   end if
-14
+13
  // LocalRAG Retrieval (at most once per round)
@@
     // Retrieve merchants, POIs, service descriptions
+16
 17
-18   end if
+18 $\mathit{information}\leftarrow\mathit{information}\cup\mathit{web\_results}\cup\mathit{rag\_results}$;
 19
-20 $\mathit{information}\leftarrow\mathit{information}\cup\mathit{web\_results}\cup\mathit{rag\_results}$;
-21
 20 if SufficientInfo($\mathit{information}$, $Q$) or $i=N$ then
 21    break;
+22
+23
 24
-25   end if
-26
-27 end for
-28
 // Phase 3: Return Results
 25 return $\mathit{information}$;
-30
-Figure 12. Prompt for Data Augmentation
-Figure 13. Prompt for Privacy Rewriting
-Figure 14. Prompt for Augmentation Validation
+26
+Algorithm 1 Algorithm of Iterative Search